### v1

# Covariant Fractons and Weitzenböck Torsion

Abstract
The relation between covariant fracton gauge theory and Møller-Hayashi-Shirafuji extension of teleparallel gravity is investigated. The former is the gauge theory of a rank two symmetric tensor with gauge symmetry given by the double derivative of a scalar parameter; the latter is a class of theories which generalises Einstein gravity, whose action is quadratic in the Weitzenböck torsion. We show that the solutions of covariant fracton gauge theory can be viewed as solutions of Møller-Hayashi-Shirafuji theory, providing a new insight in the relation between covariant fractons and gravity, and elucidating the meaning of covariant fracton theories as a new type of gauge theory.
1 Introduction and outlook
In this note we pursue the investigation on covariant fracton gauge theory and its properties, studying the relation between covariant fractons and a generalisation of teleparallel gravity.
Fracton phases, which are excitations characterised by restricted mobility due to the conservation of the dipole momentum Pretko:2017xar ; Seiberg:2020bhn ; Burnell:2021reh , were firstly introduced in condensed matter physics Vijay:2015mka , but attracted interest in a broader context. In particular, covariant fracton gauge theory, introduced for the first time in Blasi:2022mbl , is the gauge theory of a rank two symmetric tensor with gauge symmetry given by the double derivative of a scalar parameter. Covariant fracton gauge theory elucidates the main properties of ordinary fracton excitations, in the context of the higher rank extensions of Maxwell theory Bertolini:2022ijb , and it can be thought as an extension of linearised Einstein gravity, invariant with respect to longitudinal diffeomorphisms Bertolini:2023juh ; Afxonidis:2023pdq . In particular, the stability of fracton gravity Afxonidis:2024tph , three-dimensional Chern-Simons-like models with covariant fracton gauge symmetry Bertolini:2024yur , and the analogue of the electromagnetic duality in covariant fracton gauge theory Bertolini:2025jov have been recently investigated.
The most general quadratic action in the derivatives of the fracton gauge field $h_{\mu\nu}^{\text{(fr)}}$ can be written in terms of the two possible scalar contractions of the gauge invariant fracton field strength Rovere:2024nwc  [^1] It appears also in Bertolini:2024apg at the linear level. $h_{\mu\nu}^{\text{(fr)}}$ and $f_{\mu\nu\varrho}^{(\text{fr})}$ are respectively denoted with $h_{\mu\nu}$ and $f_{\mu\nu\varrho}$ in Rovere:2024nwc .
$f^{\text{(fr)}}_{\mu\nu\varrho}=\partial_{\mu}\,h_{\nu\varrho}^{\text{(fr)}}-\partial_{\nu}\,h_{\mu\varrho}^{\text{(fr)}},$ (1.1)
which is analogous to the field strength in electromagnetism. So, the action of the theory reads
$S_{\text{fr}}=\int\mathrm{d}^{d}x\,(\alpha\,{f^{\text{(fr)}}}^{\mu\nu\varrho}\,f^{\text{(fr)}}_{\mu\nu\varrho}+\beta\,{f^{\text{(fr)}}}^{\mu\nu}{}_{\nu}\,{f^{\text{(fr)}}}_{\mu\varrho}{}^{\varrho}),$ (1.2)
where $\alpha$ and $\beta$ are arbitrary constants. Taking $\alpha$ and $\beta$ such that $\alpha+\frac{1}{2}\,\beta=0$, and replacing the fracton gauge field with the parturbation of the metric in linearised gravity, one gets the Fierz-Pauli action (see (2.11)) Fierz:1939ix , which is invariant under the full group of linearised diffeomorphisms.
Since the whole dynamics of covariant fracton gauge theory is encoded in the fracton field strength, one might ask which its gravitational analogue is. We observe in this note that the covariant fracton field strength is the analogue of the linearised Weitzenböck torsion in gravity, when the latter has vanishing totally antisymmetric part. The first attempts of a torsion formulation of General Relativity, usually known as teleparallel gravity, date back to Einstein Einstein1928 , who unsuccessfully tried to unificate gravitation and electromagnetism in four dimensions, using the sixteen degrees of freedom of the vierbein, according to the works of Cartan Cartan1922 ; Cartan1923 ; Cartan1924 ; Cartan1925 and Weitzenböck Weitzenbock1928 . For the history and prehistory of the teleparallel gravity and its extension see for example Shirafuji:1995xc . In teleparallel gravity the dynamical variables of General Relativity are encoded in the torsion of a flat connection, known as Weitzenböck connection:
$W_{\mu}{}^{\varrho}{}_{\nu}=e^{\varrho}{}_{a}\,\partial_{\mu}\,e_{\nu}{}^{a}.$ (1.3)
The notion of Weitzenböck torsion is very useful for its geometrical implications, since it was employed in the geometrisation of gauge symmetries in the context of Scherk-Schwarz compactifications Scherk:1979zr , and in generalised geometry Coimbra:2011ky ; Aldazabal:2013mya ; Lee:2014mla .
Møller studied extensions of teleparallel gravity, by considering the wider class of theories defined by the most general action quadratic in the Weitzenböck torsion Moller1961a ; Moller1961b ; Moller1978 . At the same time, independently, the possible phenomenological implications of the model were firstly studied by Hayashi and Shirafuji Hayashi:1979qx . The model can be also viewed as the gauge theory of translations Hayashi:1967se . The action for the theory considered by Møller and by Hayashi and Shirafuji, which we will call Møller-Hayashi-Shirafuji theory (mhs), reads
$S_{\textsc{mhs}}=\frac{1}{2}\int\mathrm{d}^{d}x\,e\,(\alpha_{1}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\alpha_{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}+\alpha_{3}\,T^{\mu\nu}{}_{\nu}\,T_{\mu\varrho}{}^{\varrho}),$ (1.4)
where $T_{\mu}{}^{\varrho}{}_{\nu}=W_{[\mu}{}^{\varrho}{}_{\nu]}$ is the Weitzenböck torsion, which is a well-defined tensor, and $\alpha_{1},\alpha_{2},\alpha_{3}$ are arbitrary constants. We will show that all the solutions of the equations of motion of covariant fracton gauge theory are solutions of the equations of motion of linearised mhs theory.
To make the embedding more precise, we will take into account the brst formulation Becchi:1974md ; Becchi:1974xu ; Becchi:1975nq ; Tyutin:1975qk of covariant fracton gauge theory. In this setup, the gauge parameter is replaced by an anticommuting scalar field (fracton ghost) $\lambda$, and the gauge transformation is realised by a nilpotent differential operator $s$ on the space of the fracton gauge field and of the ghost and their derivatives, evaluated in the same spacetime point:
$s\,\lambda=0,$ (1.5a)
$s\,h_{\mu\nu}^{(\text{fr})}=\partial_{\mu}\,\partial_{\nu}\,\lambda.$ (1.5b)
The local $s$-cohomology, which classifies actions and anomalies of the theory, is generated by the undifferentiated ghost and by the fracton field strength and its derivatives, as in the analogous case of electromagnetism Rovere:2024nwc . We will show that the brst transformations of covariant fracton gauge theory can be conveniently formulated in an equivalent way, by adding two trivial doublets. The resulting transformations are
$s\,\mathring{\xi}_{\mu}=0,$ (1.6a)
$s\,\lambda=0,$ (1.6b)
$s\,h_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}+\partial_{\nu}\,\mathring{\xi}_{\mu},$ (1.6c)
$s\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},$ (1.6d)
$s\,r_{\mu}=\mathring{\xi}-\tfrac{1}{2}\,\partial_{\mu}\,\lambda,$ (1.6e)
$s\,B_{\mu\nu}=b_{\mu\nu}-\partial_{\mu}\,r_{\nu}+\partial_{\nu}\,r_{\mu},$ (1.6f)
where, in particular, the symmetric rank two tensor $h_{\mu\nu}$, which replaces the fracton gauge field, shifted by the symmetrised first derivative of $r_{\mu}$,
$h_{\mu\nu}=h_{\mu\nu}^{\text{(fr)}}+\partial_{\mu}\,r_{\nu}+\partial_{\nu}\,r_{\mu},$ (1.7)
enjoys the gauge transformation of linearised gravity, given by linearised diffeomorphisms and generated by the anticommuting ghost $\mathring{\xi}$. $h_{\mu\nu}$ and $b_{\mu\nu}$, the latter being an antisymmetric rank two tensor, can be respectively viewed as the symmetric and the antisymmetric part of the linearised vielbein. In Einstein gravity, in which only $h_{\mu\nu}$ is propagating, local Lorentz transformations are included in order to gauge away $b_{\mu\nu}$. In the above transformation there is no local Lorentz transformations: this is the case of Weitzenböck geometry, which is flat and it is invariant only under global Lorentz transformations Hayashi:1977jd . Nevertheless, the last of the above transformations requires $b_{\mu\nu}$ to be cohomologically equivalent to a $\mathrm{d}$-exact two form (i.e. they differ by an $s$-exact term, which vanishes in the cohomology space). As we will see, the consequence is that the linearised Weitzenböck torsion is cohomologically equivalent to (one half) the fracton field strength, and so the totally antisymmetric part of the linearised Weitzenböck torsion is cohomologically trivial. Finally, the transformation of $r_{\mu}$ serves to implement the restriction of linearised diffeomorphisms to the “longitudinal” ones, bringing to the covariant fracton gauge transformations as they were firstly introduced Blasi:2022mbl .
In Bertolini:2024apg a possible non-linear extension of covariant fracton gauge theory is discussed, by means of the Frölicher-Nijenhuis bracket, which acts on vector-valued differential forms and it coincides with the Lie derivative between vector fields when it acts on vectors FroNij1 ; FroNij2 ; FroNij3 . The non-linear extension of the fracton gauge transformation is realised by using a vector parameter, which is restricted to be the derivative of a scalar parameter only at the linear level; the gauge field is replaced by a vector-valued one-form, and a constraint is added in order to make it symmetric; a possible non-linear equation of motion is postulated, without an action from which to deduce it; the linearisation of such an equation is the equation of motion of covariant fracton gauge theory (with $\beta=0$ in (1.2)).
It would be interesting to exploit the relation between covariant fractons and Weitzenböck torsion in understanding deeper this non-linear extension. In particular, it could be nice to see if it is possible to relate the postulated non-linear equation of motion with the equations of mhs theory, since it seems natural to identity the vector-valued one form with the linearised vielbein.
The rest of the paper is organised as follows: in Section 2 covariant fracton theories are formulated in a way which makes manifest the relation with torsion formulation of gravity; in Section 3, 4, and 5 torsion formulation of General Relativity, the mhs theory, and its linearisation are reviewed; Section 6 is devoted to the embedding of covariant fractons in torsion gravity; finally, in Section 7 the brst formulation is investigated.
2 Covariant fracton gauge theory in “torsion-like formulation”
In this section covariant fracton gauge theory Blasi:2022mbl is formulated in a way in which it will be natural the embedding of in gravitational theories in torsion formulation.
Covariant fracton gauge theory is defined by a symmetric rank two tensor $h^{(\text{fr})}_{\mu\nu}$ and by a scalar gauge parameter, promoted to an anticommuting ghost $\lambda$ in brst formalism Becchi:1974md ; Becchi:1974xu ; Becchi:1975nq ; Tyutin:1975qk . The nilpotent brst transformations are
$s\,h^{(\text{fr})}_{\mu\nu}=\partial_{\mu}\,\partial_{\nu}\,\lambda,\quad s\,\lambda=0.$ (2.1)
One defines the fracton field strength
$f^{(\text{fr})}_{\mu\nu\varrho}:=\partial_{\mu}\,h^{(\text{fr})}_{\nu\varrho}-\partial_{\nu}\,h^{(\text{fr})}_{\mu\varrho},$ (2.2)
which is antisymmetric in the first two indices, and whose totally antisymmetric part vanishes (algebraic Bianchi identity). Moreover, the fracton field strength is brst invariant and it satisfies a differential Bianchi identity:
$s\,f^{\text{(fr)}}_{\mu\nu\varrho}=0,$ (2.3a)
$f^{\text{(fr)}}_{\mu\nu\varrho}+f^{\text{(fr)}}_{\nu\varrho\mu}+f^{\text{(fr)}}_{\varrho\mu\nu}=0,$ (2.3b)
$\partial_{\sigma}\,f^{\text{(fr)}}_{\mu\nu\varrho}+\partial_{\mu}\,f^{\text{(fr)}}_{\nu\sigma\varrho}+\partial_{\nu}\,f^{\text{(fr)}}_{\sigma\mu\varrho}=0.$ (2.3c)
The most general invariant quadratic action in the derivative of $h^{(\text{fr})}_{\mu\nu}$ can be written in terms of the fracton field strength, since the local brst cohomology of ghost number zero is generated by the fracton field strength and its derivatives Rovere:2024nwc . There are only two independent quadratic scalar contractions of the fracton field strength, since it transforms in the mixed Young tableau $(2,1)$. [^2] We denote with $(\lambda_{1},\dots,\lambda_{n})$ the Young tableau with $n$ rows and $\lambda_{i}$ boxes in the $i$-th row, $i=1,\dots,n$. So, the most general quadratic invariant action can be written as
$S_{\text{fr}}$ $=\int\mathrm{d}^{d}x\,(\alpha\,{f^{\text{(fr)}}}^{\mu\nu\varrho}\,f^{\text{(fr)}}_{\mu\nu\varrho}+\beta\,{f^{\text{(fr)}}}^{\mu\nu}{}_{\nu}\,{f^{\text{(fr)}}}_{\mu\varrho}{}^{\varrho})=$
$=\int\mathrm{d}^{d}x\,(2\,\alpha\,\partial^{\varrho}\,h^{\text{(fr)}\,\mu\nu}\,\partial_{\varrho}\,h^{\text{(fr)}}_{\mu\nu}-2\,(\alpha-\tfrac{\beta}{2})\,\partial^{\nu}\,h^{\text{(fr)}}_{\mu\nu}\,\partial_{\varrho}\,h^{\text{(fr)}\,\mu\varrho}\,+$
$\qquad+\beta\,\partial_{\mu}\,h^{\text{(fr)}\,\nu}{}_{\nu}\,\partial^{\mu}\,h^{\text{(fr)}\,\varrho}{}_{\varrho}-2\,\beta\,\partial_{\mu}\,h^{\text{(fr)}\,\nu}{}_{\nu}\,\partial_{\varrho}\,h^{\text{(fr)}\,\mu\varrho}),$ (2.4)
for arbitrary constants $\alpha$ and $\beta$. The corresponding equations of motion require the following symmetric tensor to vanish
$E^{\text{(fr)}}_{\mu\nu}:=-\frac{1}{2}\,\frac{\delta\,S_{\text{fr}}}{\delta\,{h^{(\text{fr})}}^{\mu\nu}}=\alpha\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho}{}_{(\mu\nu)}-\tfrac{\beta}{2}\,\partial_{(\mu}\,{f^{\text{(fr)}}}_{\nu)\varrho}{}^{\varrho}+\beta\,\eta_{\mu\nu}\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho\sigma}{}_{\sigma},$ (2.5)
where the round brackets among indices denote symmetrisation without any numerical coefficient. [^3] For example, $A_{(\mu\nu)}=A_{\mu\nu}+A_{\nu\mu}$. Similarly, $A_{[\mu\nu]}=A_{\mu\nu}-A_{\nu\mu}$. It is useful to define the following tensor
$R^{\text{(fr)}}_{\mu\nu\varrho\sigma}$ $=\tfrac{1}{2}\,\partial_{\sigma}\,\partial_{\mu}\,h^{(\text{fr})}_{\nu\varrho}-\tfrac{1}{2}\,\partial_{\sigma}\,\partial_{\nu}\,h^{(\text{fr})}_{\mu\varrho}-\tfrac{1}{2}\,\partial_{\varrho}\,\partial_{\mu}\,h^{(\text{fr})}_{\nu\sigma}+\tfrac{1}{2}\,\partial_{\varrho}\,\partial_{\nu}\,h^{(\text{fr})}_{\mu\sigma}=$
$=\tfrac{1}{2}\,\partial_{\sigma}\,f^{\text{(fr)}}_{\mu\nu\varrho}-\tfrac{1}{2}\,\partial_{\varrho}\,f^{\text{(fr)}}_{\mu\nu\sigma},$ (2.6)
which enjoys the same symmetries as the Riemann tensor, being antisymmetric in $\mu\nu$ and in $\varrho\sigma$ and symmetric in $\mu\nu\leftrightarrow\varrho\sigma$. The corresponding “Ricci tensor” and “Ricci scalar” respectively read [^4] Consider that $\partial_{[\mu}\,{f^{\text{(fr)}}}_{\nu]\varrho}{}^{\varrho}-\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho}{}_{[\mu\nu]}=0$ as a consequence of (2.3b)–(2.3c).
$R^{(\text{fr})}_{\mu\nu}$ $={R^{\text{(fr)}}}^{\varrho}{}_{\mu\varrho\nu}=-\tfrac{1}{4}\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho}{}_{(\mu\nu)}-\tfrac{1}{4}\,\partial_{(\mu}\,{f^{\text{(fr)}}}_{\nu)\varrho}{}^{\varrho},$ (2.7)
$R^{(\text{fr})}$ $={R^{(\text{fr})}}_{\mu}{}^{\mu}=-\partial_{\mu}\,{f^{\text{(fr)}}}^{\mu\nu}{}_{\nu}.$ (2.8)
Writing $\partial_{(\mu}\,{f^{\text{(fr)}}}_{\nu)\varrho}{}^{\varrho}$ and $\partial_{\mu}\,{f^{\text{(fr)}}}^{\mu\nu}{}_{\nu}$ in terms of the “Ricci tensor” and of the “Ricci scalar” in the equations of motion (2.5), one recognises the analogue of the linearised Einstein tensor:
$E^{\text{(fr)}}_{\mu\nu}=2\,\beta\,(R^{\text{(fr)}}_{\mu\nu}-\tfrac{1}{2}\,R^{\text{(fr)}}\,\eta_{\mu\nu})+(\alpha+\tfrac{\beta}{2})\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho}{}_{(\mu\nu)}.$ (2.9)
Indeed, if one replace $h^{\text{(fr)}}_{\mu\nu}$ in the action (2) with a rank two symmetric tensor $h_{\mu\nu}$, thought as the perturbation of a metric $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$, and $\alpha$ and $\beta$ are such that
$\alpha+\tfrac{\beta}{2}=0,$ (2.10)
choosing the normalisation $\alpha=\frac{1}{2}$, one recovers the Fierz-Pauli action Fierz:1939ix
$S_{\textsc{fp}}=\tfrac{1}{2}\int\mathrm{d}^{d}x\,(\partial_{\varrho}\,h_{\mu\nu}\,\partial^{\varrho}\,h^{\mu\nu}-2\,\partial^{\nu}\,h_{\mu\nu}\,\partial_{\varrho}\,h^{\mu\varrho}+2\,\partial_{\mu}\,h_{\nu}{}^{\nu}\,\partial_{\varrho}\,h^{\mu\varrho}-\partial_{\mu}\,h_{\nu}{}^{\nu}\,\partial^{\mu}\,h_{\varrho}{}^{\varrho}).$ (2.11)
Consistently, imposing the condition (2.10) in the equations of motion (2.9), the last term vanishes and one gets the linearised Einstein equations.
Using the equations of motion in the form (2.9), it is manifest that the first divergence is not identically zero, unless in the Fierz-Pauli case, since only the analogue of the Einstein tensor is transverse:
$\partial^{\mu}\,E^{\text{(fr)}}_{\mu\nu}=(\alpha+\tfrac{\beta}{2})\,\partial^{\mu}\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho}{}_{(\mu\nu)}.$ (2.12)
Instead, one gets a transversality condition for all values of $\alpha$ and $\beta$ by taking the second divergence, as a consequence of the invariance under the fracton gauge transformations (2.1),
$\partial^{\mu}\,\partial^{\nu}\,E^{\text{(fr)}}_{\mu\nu}=(\alpha+\tfrac{\beta}{2})\,\partial^{\mu}\,\partial^{\nu}\,\partial^{\varrho}\,f^{\text{(fr)}}_{\mu\nu\varrho}=0,$ (2.13)
which vanishes since the fracton field strength is antisymmetric in its first two indices.
3 Review of teleparallel gravity
In the usual formulation of General Relativity, one chooses a connection $\Gamma_{\mu}{}^{\varrho}{}_{\nu}$ compatible with the metric $g_{\mu\nu}$, such that the torsion vanishes and the curvature encodes the dynamical variables:
$\ignorespaces\nabla_{\varrho}(\Gamma)\,g_{\mu\nu}$ $=\partial_{\varrho}\,g_{\mu\nu}-\Gamma_{\varrho}{}^{\sigma}{}_{\mu}\,g_{\sigma\nu}-\Gamma_{\varrho}{}^{\sigma}{}_{\nu}\,g_{\mu\sigma}=0,$ (3.1a)
$T_{\mu}{}^{\varrho}{}_{\nu}(\Gamma)$ $=\Gamma_{\mu}{}^{\varrho}{}_{\nu}-\Gamma_{\nu}{}^{\varrho}{}_{\mu}=0,$ (3.1b)
$R_{\mu\nu}{}^{\varrho}{}_{\sigma}(\Gamma)$ $=\partial_{\mu}\,\Gamma_{\nu}{}^{\varrho}{}_{\sigma}-\partial_{\nu}\,\Gamma_{\mu}{}^{\varrho}{}_{\sigma}+\Gamma_{\mu}{}^{\varrho}{}_{\tau}\,\Gamma_{\nu}{}^{\tau}{}_{\sigma}-\Gamma_{\nu}{}^{\varrho}{}_{\tau}\,\Gamma_{\mu}{}^{\tau}{}_{\sigma}.$ (3.1c)
As known, there is a unique choice for such a connection, the Levi-Civita connection:
$\Gamma_{\mu}{}^{\varrho}{}_{\nu}=\tfrac{1}{2}\,g^{\varrho\sigma}\,(\partial_{\mu}\,g_{\nu\sigma}+\partial_{\nu}\,g_{\mu\sigma}-\partial_{\sigma}\,g_{\mu\nu}).$ (3.2)
In the following we will denote with $\nabla_{\mu}$ the covariant derivative with respect to the Levi-Civita connection $\Gamma_{\mu}{}^{\varrho}{}_{\nu}$, and with $R^{\varrho}{}_{\sigma\mu\nu}$ the Riemann tensor.
There is an alternative but equivalent formulation of General Relativity (at least classically and without boundary terms), known as teleparallel gravity, in which one chooses a connection $W_{\mu}{}^{\varrho}{}_{\nu}$, compatible with the metric, such that the dynamical variables are encoded in the torsion and the curvature vanishes Einstein1928 ; Weitzenbock1928 :
$\ignorespaces\nabla_{\varrho}(W)\,g_{\mu\nu}$ $=\partial_{\varrho}\,g_{\mu\nu}-W_{\varrho}{}^{\sigma}{}_{\mu}\,g_{\sigma\nu}-W_{\varrho}{}^{\sigma}{}_{\nu}\,g_{\mu\sigma}=0,$ (3.3a)
$T_{\mu}{}^{\varrho}{}_{\nu}(W)$ $=W_{\mu}{}^{\varrho}{}_{\nu}-W_{\nu}{}^{\varrho}{}_{\mu},$ (3.3b)
$R_{\mu\nu}{}^{\varrho}{}_{\sigma}(W)$ $=\partial_{\mu}\,W_{\nu}{}^{\varrho}{}_{\sigma}-\partial_{\nu}\,W_{\mu}{}^{\varrho}{}_{\sigma}+W_{\mu}{}^{\varrho}{}_{\tau}\,W_{\nu}{}^{\tau}{}_{\sigma}-W_{\nu}{}^{\varrho}{}_{\tau}\,W_{\mu}{}^{\tau}{}_{\sigma}=0.$ (3.3c)
While in the usual formulation of General Relativity the gravitational interaction is due to the curvature of spacetime, in teleparallel gravity the gravitational interaction is mediated by a Lorentz-like force in a flat spacetime, as in Yang-Mills theories deAndrade:1997gka . Indeed, while in the former case the action is linear in the curvature, in the latter one the action is quadratic in the torsion.
Since the spacetime manifold is supposed to be flat, then it is possible to choose a system of locally inertial coordinates valid everywhere. In other words, the frame or inverse vielbein $e^{\mu}{}_{a}$ is a set of $d$ vectors in $d$ dimensions in the tangent space of the manifold such that, when it is evaluated in a point, it forms a basis for the tangent space in that point, for all the points of the manifold. In such a case the manifold is said parallelisable, and it is possible to choose a connection such that the vielbein is parallel or covariantly constant
$\partial_{\mu}\,e_{\nu}{}^{a}-W_{\mu}{}^{\varrho}{}_{\nu}\,e_{\varrho}{}^{a}=0,$ (3.4)
The unique connection solving the previous condition is the Weitzenböck connection Weitzenbock1928
$W_{\mu}{}^{\varrho}{}_{\nu}=e^{\varrho}{}_{a}\,\partial_{\mu}\,e_{\nu}{}^{a},$ (3.5)
which consistently has vanishing curvature, since it is the Cartan-Maurer form out of the vielbein. [^5] The spin connection corresponding to the Weitzenböck connection is zero (Weitzenböck condition). See A for more details. Moreover, the condition (3.4) implies the compatibility condition (3.3a), since
$g_{\mu\nu}=e_{\mu}{}^{a}\,e_{\nu}{}^{b}\,\eta_{ab}.$ (3.6)
In the following we will denote with $D_{\mu}$ the covariant derivative with respect to the Weitzenböck connection $W_{\mu}{}^{\varrho}{}_{\nu}$, and with $T_{\mu}{}^{\varrho}{}_{\nu}$ the Weitzenböck torsion,
$T_{\mu}{}^{\varrho}{}_{\nu}=e^{\varrho}{}_{a}\,\partial_{[\mu}\,e_{\nu]}{}^{a},$ (3.7)
antisymmetric in the first and in the last index.
The Weitzenböck torsion satisfies the following Bianchi identity
$D_{[\mu}\,T_{\nu]}{}^{\sigma}{}_{\varrho}+D_{\varrho}\,T_{\mu}{}^{\sigma}{}_{\nu}-T_{[\mu}{}^{\sigma\tau}\,T_{\nu]\tau\varrho}-T_{\mu\tau\nu}\,T_{\varrho}{}^{\sigma\tau}=0,$ (3.8)
whose trace, which will be useful in the following, reads
$D_{[\mu}\,T_{\nu]\varrho}{}^{\varrho}+D_{\varrho}\,T_{\mu}{}^{\varrho}{}_{\nu}+T_{\mu\varrho\nu}\,T^{\varrho\sigma}{}_{\sigma}=0.$ (3.9)
Using the derivative of the vielbein determinant $\partial_{\mu}\,e=e\,W_{\mu}{}^{\varrho}{}_{\varrho}$, one can show that the following identity holds for the Weitzenböck covariant divergence, for any vector $V^{\mu}$,
$e\,D_{\mu}\,V^{\mu}=\partial_{\mu}\,(e\,V^{\mu})+e\,T_{\mu}{}^{\mu}{}_{\nu}\,V^{\nu}.$ (3.10)
The relation between the Levi-Civita connection and the Weitzenböck is given by:
$\Gamma_{\mu}{}^{\varrho}{}_{\nu}=W_{\mu}{}^{\varrho}{}_{\nu}+K_{\mu}{}^{\varrho}{}_{\nu}=W_{\mu}{}^{\varrho}{}_{\nu}+\tfrac{1}{2}\,(T_{\mu\nu}{}^{\varrho}+T_{\nu\mu}{}^{\varrho}-T_{\mu}{}^{\varrho}{}_{\nu}),$ (3.11)
where the tensor $-K_{\mu}{}^{\varrho}{}_{\nu}$ is the contorsion. Using the vanishing of the curvature of the Weitzenböck connection and the relation (3.11), one can write the Riemann tensor in terms of the contorsion in the following way:
$R^{\mu}{}_{\nu\varrho\sigma}=D_{\varrho}\,K_{\sigma}{}^{\mu}{}_{\nu}-D_{\sigma}\,K_{\rho}{}^{\mu}{}_{\nu}+T_{\varrho}{}^{\tau}{}_{\sigma}\,K_{\tau}{}^{\mu}{}_{\nu}+K_{\varrho}{}^{\mu}{}_{\tau}\,K_{\sigma}{}^{\tau}{}_{\nu}-K_{\sigma}{}^{\mu}{}_{\tau}\,K_{\varrho}{}^{\tau}{}_{\nu}.$ (3.12)
Therefore, using the trace of the Bianchi identity (3.9), the Ricci tensor is
$R_{\mu\nu}=R^{\varrho}{}_{\mu\varrho\nu}$ $=\tfrac{1}{2}\,D_{\varrho}\,T_{(\mu\nu)}{}^{\varrho}-\tfrac{1}{2}\,D_{(\mu}\,T_{\nu)\varrho}{}^{\varrho}+\tfrac{1}{2}\,T_{(\mu\nu)\varrho}\,T^{\varrho\sigma}{}_{\sigma}\,+$
$+\tfrac{1}{4}\,T_{\varrho\mu\sigma}\,T^{\varrho}{}_{\nu}{}^{\sigma}-\tfrac{1}{2}\,T_{\mu\varrho\sigma}\,T_{\nu}{}^{\varrho\sigma}-\tfrac{1}{2}\,T_{\mu\varrho\sigma}\,T_{\nu}{}^{\sigma\varrho},$ (3.13)
whose the Ricci scalar is
$R=R^{\mu}{}_{\mu}=-2\,D_{\mu}\,T^{\mu\nu}{}_{\nu}-\tfrac{1}{4}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}-\tfrac{1}{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}-T_{\mu\nu}{}^{\nu}\,T^{\mu\varrho}{}_{\varrho}.$ (3.14)
Finally, using the Weitzenböck covariant divergence (3.10), one gets the following identity
$e\,R=-\partial_{\mu}\,(2\,e\,T^{\mu\nu}{}_{\nu})-e\,(\tfrac{1}{4}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\tfrac{1}{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}-T_{\mu\nu}{}^{\nu}\,T^{\mu\varrho}{}_{\varrho}),$ (3.15)
which shows that, up to a boundary term, the Einstein-Hilbert action admits an equivalent formulation in terms of a linear combination of the quadratic scalar contractions of the Weitzenböck torsion Moller1961a ; Moller1961b ; Moller1978 ; Hayashi:1967se ; Hayashi:1979qx ; deAndrade:1997gka :
$S_{\textsc{eh}}=-\frac{1}{2}\int\mathrm{d}^{d}x\,e\,R=\frac{1}{2}\int\mathrm{d}^{d}x\,e\,\mathopen{}\mathclose{{}\left(\frac{1}{4}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\frac{1}{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}-T_{\mu\nu}{}^{\nu}\,T^{\mu\varrho}{}_{\varrho}}\right),$ (3.16)
where we set $\kappa=8\,\pi\,G=1$.
4 Review of Møller-Hayashi-Shirafuji theory
The relation (3.16) shows that the Einstein-Hilbert action consists in a linear combination of three quadratic scalar contractions of the Weitzenböck torsion with precise relative coefficients. In Moller1961a ; Moller1961b ; Moller1978 ; Hayashi:1967se ; Hayashi:1979qx the theory defined by the same action with arbitrary coefficients is considered:
$S_{\textsc{mhs}}=\frac{1}{2}\int\mathrm{d}^{d}x\,e\,(\alpha_{1}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\alpha_{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}+\alpha_{3}\,T^{\mu\nu}{}_{\nu}\,T_{\mu\varrho}{}^{\varrho}).$ (4.1)
We will call the theory described by this action Møller-Hayashi-Shirafuji theory (mhs). [^6] Hayashi and Shirafuji called it New General Relativity Hayashi:1979qx . See Appendix C for a comparison with definitions and conventions in this paper and in Moller1978 and in Hayashi:1979qx .
The peculiar combination picking the case of General Relativity, which is
$\alpha_{1}=\tfrac{1}{4},\quad\alpha_{2}=\tfrac{1}{2},\quad\alpha_{3}=-1,$ (4.2)
is captured by formally requiring the local Lorentz invariance, which indeed is the necessary requirement in the vielbein formulation of General Relativity, beside diffeomorphism invariance, in order to recover the right number of degrees of freedom Cho:1975dh .
The mhs action is the most general quadratic action in the Weitzenböck torsion, since there are only three possible independent quadratic scalar contractions of the torsion. Indeed, the torsion has three irreducible components:
$T_{\mu\varrho\nu}\rightarrow(1,1)\otimes(1)=(1,1,1)\otimes(2,1)_{t}\otimes(1),$ (4.3)
where $(\lambda_{1},\lambda_{2},\dots,\lambda_{p})$ denotes a Young tableau with $p$ rows and $\lambda_{i}$ boxes in the $i$-th row, $i=1,\dots,p$; and $t$ denotes the traceless part. $(1,1,1)$ is the totally antisymmetric part $T_{[\mu\nu\varrho]}$, and $(1)$ is the unique trace $T_{\mu\nu}{}^{\nu}$. Namely, the contraction out of two traces is $T_{\mu\nu}{}^{\nu}\,T^{\mu\varrho}{}_{\varrho}$; the contraction out of two totally antisymmetric tensors is $T^{[\mu\nu\varrho]}\,T_{[\mu\nu\varrho]}$, which is proportional to $T^{\mu\nu\varrho}\,T_{[\mu\nu\varrho]}$; the scalar contraction involving the remaining irreducible component can be replaced by $T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}$; and the second contraction can be replaced by $T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}$, since
$T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}=\tfrac{1}{2}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}-\tfrac{1}{4}\,T^{\mu\nu\varrho}\,T_{[\mu\nu\varrho]}.$ (4.4)
Therefore, the three contractions in (4.1) can be chosen as an independent basis.
Using the variations
$\delta\,e=-e\,e_{\mu}{}^{a}\,\delta\,e^{\mu}{}_{a},\quad\delta\,T_{\mu}{}^{\varrho}{}_{\nu}=e_{[\mu}{}^{a}\,D_{\nu]}\,\delta^{\varrho}{}_{a},$ (4.5)
one can compute the equations of motion of the tree pieces in (4.1) Moller1978 :
$\ignorespaces\frac{e_{\mu a}}{e}\,\frac{\delta}{\delta e^{\nu}{}_{a}}\,\int\mathrm{d}^{d}x\,e\,T^{\alpha\beta\gamma}\,T_{\alpha\beta\gamma}=$ $\,-g_{\mu\nu}\,T^{\alpha\beta\gamma}\,T_{\alpha\beta\gamma}-2\,T_{\alpha\mu\beta}\,T^{\alpha}{}_{\nu}{}^{\beta}-4\,T_{\mu\nu}{}^{\alpha}\,T_{\alpha\beta}{}^{\beta}\,+$
$+4\,T_{\mu\alpha\beta}\,T_{\nu}{}^{\alpha\beta}-4\,D_{\alpha}\,T_{\mu\nu}{}^{\alpha},$ (4.6a)
$\frac{e_{\mu a}}{e}\,\frac{\delta}{\delta e^{\nu}{}_{a}}\,\int\mathrm{d}^{d}x\,e\,T^{\alpha\beta\gamma}\,T_{\alpha\gamma\beta}=$ $\,-g_{\mu\nu}\,T^{\alpha\beta\gamma}\,T_{\alpha\gamma\beta}-2\,T_{\mu\alpha\nu}\,T^{\alpha\beta}{}_{\beta}+2\,T_{\mu\alpha\beta}\,T_{\nu}{}^{\beta\alpha}\,+$
$-2\,T_{\nu\mu\alpha}\,T^{\alpha\beta}{}_{\beta}-2\,D_{\alpha}\,T_{\mu}{}^{\alpha}{}_{\nu}-2\,D_{\alpha}\,T_{\nu\mu}{}^{\alpha},$ (4.6b)
$\frac{e_{\mu a}}{e}\,\frac{\delta}{\delta e^{\nu}{}_{a}}\,\int\mathrm{d}^{d}x\,e\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}=$ $\,g_{\mu\nu}\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}+2\,g_{\mu\nu}\,D_{\alpha}\,T^{\alpha\beta}{}_{\beta}-2\,D_{\nu}\,T_{\mu\alpha}{}^{\alpha}.$ (4.6c)
Therefore, the equations of motion $E_{\mu\nu}:=\frac{e_{\mu a}}{e}\frac{\delta S_{\textsc{mhs}}}{\delta e^{\nu}{}_{a}}=0$, decomposed in symmetric and antisymmetric part $E_{\mu\nu}=E_{\mu\nu}^{(\text{sym})}+E_{\mu\nu}^{\text{(anti)}}$, with $E_{\mu\nu}^{\text{(sym)}}=E_{\nu\mu}^{\text{(sym)}}$, and $E_{\mu\nu}^{\text{(anti)}}=-E_{\nu\mu}^{\text{(anti)}}$, respectively read
$\ignorespaces E_{\mu\nu}^{\text{(sym)}}=$  $\,2\,\alpha_{1}\,T_{\mu\alpha\beta}\,T_{\nu}{}^{\alpha\beta}+\alpha_{2}\,T_{\mu\alpha\beta}\,T_{\nu}{}^{\beta\alpha}-\alpha_{1}\,T_{\alpha\mu\beta}\,T^{\alpha}{}_{\nu}{}^{\beta}\,+$
$-(\alpha_{1}+\tfrac{\alpha_{2}}{2})(T_{(\mu\nu)}{}^{\alpha}\,T_{\alpha\beta}{}^{\beta}+D_{\alpha}\,T_{(\mu\nu)}{}^{\alpha})-\tfrac{\alpha_{3}}{2}\,D_{(\mu}\,T_{\nu)\alpha}{}^{\alpha}\,+$
$-\tfrac{1}{2}\,g_{\mu\nu}\,(\alpha_{1}\,T^{\alpha\beta\gamma}\,T_{\alpha\beta\gamma}+\alpha_{2}\,T^{\alpha\beta\gamma}\,T_{\alpha\gamma\beta}-\alpha_{3}\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}-2\,\alpha_{3}\,D_{\alpha}\,T^{\alpha\beta}{}_{\beta}),$ (4.7a)
$E_{\mu\nu}^{\text{(anti)}}=$ $\,(\tfrac{\alpha_{2}}{2}-\alpha_{1})\,(T_{[\mu\nu]}{}^{\alpha}\,T_{\alpha\beta}{}^{\beta}+D_{\alpha}\,T_{[\mu\nu]}{}^{\alpha})-(\alpha_{2}+\tfrac{\alpha_{3}}{2})\,(T_{\mu\alpha\nu}\,T^{\alpha\beta}{}_{\beta}+D_{\alpha}\,T_{\mu}{}^{\alpha}{}_{\nu}).$ (4.7b)
where in the antisymmetric part we use the trace of the Bianchi identity (3.9). Using the choice (4.2), the antisymmetric equation is trivial (and this is the unique case) and the symmetric part, using (3.13) and (3.14), is recognised to be the Einstein equations. To make this manifest, use (3.13) and (3.14) to replace $D_{(\mu}\,T_{\nu)\alpha}{}^{\alpha}$ and $D_{\alpha}\,T^{\alpha\beta}{}_{\beta}$:
$D_{(\mu}\,T_{\nu)\alpha}{}^{\alpha}=$ $\,-2\,R_{\mu\nu}-T_{\mu\alpha\beta}\,T_{\nu}{}^{\alpha\beta}-T_{\mu\alpha\beta}\,T_{\nu}{}^{\beta\alpha}\,+$
$+\tfrac{1}{2}\,T_{\alpha\mu\beta}\,T^{\alpha}{}_{\nu}{}^{\beta}+T_{(\mu\nu)}{}^{\alpha}\,T_{\alpha\beta}{}^{\beta}+D_{\alpha}\,T_{(\mu\nu)}{}^{\alpha},$ (4.8)
$D_{\alpha}\,T^{\alpha\beta}{}_{\beta}=$ $\,-\tfrac{1}{2}\,R-\tfrac{1}{8}\,T^{\alpha\beta\gamma}\,T_{\alpha\beta\gamma}-\tfrac{1}{4}\,T^{\alpha\beta\gamma}\,T_{\alpha\gamma\beta}-\tfrac{1}{2}\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}.$ (4.9)
Moreover, it will be useful for the following to decompose the torsion in the components in transforming in the Young tableaux $(2,1)$ and $(1,1,1)$, the first one including the trace:
$T_{\mu\nu\varrho}=f_{\mu\varrho\nu}+t_{\mu\nu\varrho},$ (4.10)
where [^7] The inverse relations read: $f_{\mu\nu\varrho}=\tfrac{1}{3}\,(T_{\mu\nu\varrho}+T_{\varrho\mu\nu}-2\,T_{\nu\varrho\mu})$ and $t_{\mu\nu\varrho}=\tfrac{1}{3}\,(T_{\mu\nu\varrho}+T_{\varrho\mu\nu}+T_{\nu\varrho\mu})$.
$f_{\mu\nu\varrho}=-f_{\nu\mu\varrho},\quad f_{[\mu\nu\varrho]}=0,\quad t_{\mu\nu\varrho}=\tfrac{1}{6}\,t_{[\mu\nu\varrho]}.$ (4.11)
These symmetries imply that the following identities are satisfied:
$\ignorespaces f_{\alpha\beta\gamma}\,f^{\alpha\gamma\beta}-\tfrac{1}{2}\,f_{\alpha\beta\gamma}\,f^{\alpha\beta\gamma}=0,$ (4.12a)
$f^{\alpha\beta}{}_{\beta}\,f_{\alpha\mu\nu}-f^{\alpha\beta}{}_{\beta}\,f_{\mu\nu\alpha}=0,$ (4.12b)
$f_{\alpha\beta\gamma}\,t^{\alpha\beta\gamma}=0,$ (4.12c)
$f_{(\mu}{}^{\alpha\beta}\,t_{\nu)\alpha\beta}+\tfrac{1}{2}\,f^{\alpha\beta}{}_{(\mu}\,t_{\nu)\alpha\beta}=0,$ (4.12d)
$-\tfrac{1}{2}\,f_{\alpha\beta\mu}\,f^{\alpha\beta}{}_{\nu}+f_{\mu\alpha\beta}\,f_{\nu}{}^{\alpha\beta}-f_{\mu\alpha\beta}\,f_{\nu}{}^{\beta\alpha}=0,$ (4.12e)
$f_{\alpha\beta\mu}\,f^{\alpha\beta}{}_{\nu}+f_{\alpha\beta(\mu}\,f_{\nu)}{}^{\alpha\beta}=0.$ (4.12f)
Using the replacement rules (4.8)–(4.9), the decomposition of the torsion (4.10), and the identities (4.12a)–(4.12f) in the equations of motion (4.7a) and (4.7b), one obtains
$E_{\mu\nu}^{\text{(sym)}}=$ $\,\alpha_{3}\,(R_{\mu\nu}-\tfrac{1}{2}\,R\,g_{\mu\nu})-(4\,\alpha_{1}-\alpha_{2}+\tfrac{\alpha_{3}}{2})\,f_{(\mu}{}^{\alpha\beta}\,t_{\nu)\alpha\beta}\,+$
$+(\alpha_{1}-\alpha_{2}-\tfrac{\alpha_{3}}{4})\,(t_{\mu\alpha\beta}\,t_{\nu}{}^{\alpha\beta}-\tfrac{1}{2}\,g_{\mu\nu}\,t_{\alpha\beta\gamma}\,t^{\alpha\beta\gamma})\,+$
$+(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,(\tfrac{1}{2}\,D_{\alpha}\,f^{\alpha}{}_{(\mu\nu)}+\tfrac{1}{2}\,f^{\alpha\beta}{}_{\beta}\,f_{\alpha(\mu\nu)}\,+$
$+f_{\mu\alpha\beta}\,f_{\nu}{}^{\beta\alpha}-\tfrac{1}{4}\,g_{\mu\nu}\,f_{\alpha\beta\gamma}\,f^{\alpha\beta\gamma}),$ (4.13)
$E_{\mu\nu}^{\text{(anti)}}=$ $\,\tfrac{1}{2}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,(D_{\alpha}\,f^{\alpha}{}_{[\mu\nu]}+f^{\alpha\beta}{}_{\beta}\,f_{\alpha[\mu\nu]}\,+$
$+(-2\,\alpha_{1}+2\,\alpha_{2}+\tfrac{\alpha_{3}}{2})\,(D_{\alpha}\,t^{\alpha}{}_{\mu\nu}+f^{\alpha\beta}{}_{\beta}\,t_{\alpha\mu\nu}).$ (4.14)
Using the choice (4.2), all the combinations of the constants $\alpha_{1},\alpha_{2},\alpha_{3}$ vanish, and the symmetric part gives the Einstein equations.
5 Linearisation of Møller-Hayashi-Shirafuji theory
The linearisation of the Møller-Hayashi-Shirafuji theory Moller1978 ; Hayashi:1979qx is obtained by defining
$e_{\mu}{}^{a}=\delta_{\mu}^{a}+\varepsilon\,\overline{h}_{\mu}{}^{a},\quad g_{\mu\nu}=\eta_{\mu\nu}+\varepsilon\,h_{\mu\nu},$ (5.1)
where $\varepsilon$ is infinitesimal. Moreover, defining
$\overline{h}_{\mu}{}^{\nu}=\overline{h}_{\mu}{}^{a}\,\delta_{a}^{\nu},\quad\overline{h}_{\mu\nu}=\overline{h}_{\mu}{}^{\varrho}\,\eta_{\varrho\nu}=\overline{h}_{\mu}{}^{a}\,\delta_{a}{}^{\varrho}\,\eta_{\varrho\nu}$ (5.2)
and combining the linearisation of $g_{\mu\nu}$ and $e_{\mu}{}^{a}$, we get that $h_{\mu\nu}$ is the symmetric part $\overline{h}_{\mu\nu}+\overline{h}_{\nu\mu}$. So, denoting with $b_{\mu\nu}=\overline{h}_{\mu\nu}-\overline{h}_{\nu\mu}$ the antisymmetric part, we can decompose
$\overline{h}_{\mu\nu}=\tfrac{1}{2}\,(h_{\mu\nu}+b_{\mu\nu}).$ (5.3)
Similarly for the inverse vielbein:
$e^{\mu}{}_{a}=\delta^{\mu}_{a}-\varepsilon\,\overline{h}_{a}{}^{\mu},\quad\overline{h}_{a}{}^{\nu}\,\delta^{a}_{\mu}=\overline{h}_{\mu}{}^{a}\,\delta_{a}^{\nu},\quad\overline{h}_{b}{}^{\mu}\,\delta^{a}_{\nu}=\overline{h}_{\nu}{}^{a}\,\delta_{b}^{\mu},$ (5.4)
which are fixed in order to recover $e_{\mu}{}^{a}\,e^{\nu}{}_{a}=\delta_{\mu}^{\nu}$ and $e_{\mu}{}^{a}\,e^{\mu}{}_{b}=\delta^{a}_{b}$.
The linearised equations of motion are
$\mathring{E}_{\mu\nu}^{\text{(sym)}}=$ $\,\alpha_{3}\,(\mathring{R}_{\mu\nu}-\tfrac{1}{2}\,\mathring{R}\,\eta_{\mu\nu})+\tfrac{1}{2}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\partial_{\alpha}\,\mathring{f}^{\alpha}{}_{(\mu\nu)},$ (5.5)
$\mathring{E}_{\mu\nu}^{\text{(anti)}}=$ $\,(-2\,\alpha_{1}+2\,\alpha_{2}+\tfrac{\alpha_{3}}{2})\,\partial_{\alpha}\,\mathring{t}^{\alpha}{}_{\mu\nu}+\tfrac{1}{2}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\partial_{\alpha}\,\mathring{f}^{\alpha}{}_{[\mu\nu]}.$ (5.6)
where
$\mathring{T}_{\mu\nu\varrho}$ $=\mathring{f}_{\mu\varrho\nu}+\mathring{t}_{\mu\nu\varrho}=\partial_{\mu}\,\overline{h}_{\varrho\nu}-\partial_{\varrho}\,\overline{h}_{\varrho\nu}=$ (5.7)
$=\tfrac{1}{2}\,(\partial_{\mu}\,h_{\varrho\nu}-\partial_{\varrho}\,h_{\mu\nu}+\partial_{\mu}\,b_{\varrho\nu}-\partial_{\varrho}\,b_{\mu\nu}),$ (5.8)
$\mathring{f}_{\mu\nu\varrho}$ $=\tfrac{1}{2}\,\partial_{[\mu}\,h_{\nu]\varrho}+\tfrac{1}{6}\,(\partial_{[\mu}\,b_{\nu]\varrho}+2\,\partial_{\varrho}\,b_{\nu\mu}),$ (5.9)
$\mathring{t}_{\mu\nu\varrho}$ $=\tfrac{1}{6}\,\partial_{[\mu}\,b_{\varrho\nu]},$ (5.10)
$\mathring{R}_{\mu\nu}$ $=\tfrac{1}{2}\,\partial_{\varrho}\,\mathring{T}_{(\mu\nu)}{}^{\varrho}-\tfrac{1}{2}\,\partial_{[\mu}\,\mathring{T}_{\nu)\varrho}{}^{\varrho}=-\tfrac{1}{2}\,\partial_{\varrho}\,\mathring{f}^{\varrho}{}_{(\mu\nu)}-\tfrac{1}{2}\,\partial_{(\mu}\,\mathring{f}_{\nu)\varrho}{}^{\varrho}=$
$=-\tfrac{1}{2}\,\partial^{2}\,h_{\mu\nu}+\tfrac{1}{2}\,\partial^{\alpha}\,\partial_{(\mu}\,h_{\nu)\alpha}-\tfrac{1}{2}\,\partial_{\mu}\,\partial_{\nu}\,h_{\alpha}{}^{\alpha},$ (5.11)
$\mathring{R}$ $=-2\,\partial_{\mu}\,\mathring{T}^{\mu\nu}{}_{\nu}=-2\,\partial_{\mu}\,\mathring{f}^{\mu\nu}{}_{\nu}=\partial_{\mu}\,\partial_{\nu}\,h^{\mu\nu}-\partial^{2}\,h_{\mu}{}^{\mu},$ (5.12)
Since the mhs theory is formulated in a quadratic way, unlike General Relativity in conventional formulation, the linearised equations of motion can be obtained directly by varying the action (4.1) with the torsion $T_{\mu\nu\varrho}$ replaced by the linearised one $\mathring{T}_{\mu\nu\varrho}$:
$\frac{1}{2}\int\mathrm{d}^{d}\,x\,(\alpha_{1}\,\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\nu\varrho}+\alpha_{2}\,\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\varrho\nu}+\alpha_{3}\,\mathring{T}^{\mu\nu}{}_{\nu}\,\mathring{T}_{\mu\varrho}{}^{\varrho}).$ (5.13)
This action can be obtained working directly at the linear level in the following way. One has to look for the most general quadratic action in the derivatives of the perturbation $\overline{h}_{\mu\nu}$, invariant under linearised diffeomorphisms, which are encoded in the following brst rules: [^8] They are the linearisation of the brst rule for the vielbein encoding the diffeomorphisms, which are generated by the Lie derivative: $s\,e_{\mu}{}^{a}=\mathcal{L}_{\xi}\,e_{\mu}{}^{a}=\xi^{\nu}\,\partial_{\nu}\,e_{\mu}{}^{a}+e_{\nu}{}^{a}\,\partial_{\mu}\,\xi^{\nu}$, $s\,\xi^{\mu}=\frac{1}{2}\,\mathcal{L}_{\xi}\,\xi^{\mu}=\xi^{\nu}\,\partial_{\nu}\,\xi^{\mu}$, setting $\xi^{\mu}=\varepsilon\,\mathring{\xi}^{\mu}$.
$s\,\overline{h}_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu},\quad s\,\mathring{\xi}^{\mu}=0.$ (5.14)
or equivalently, using (5.3),
$s\,h_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}+\partial_{\nu}\,\mathring{\xi}_{\mu},\quad s\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},\quad\quad s\,\mathring{\xi}^{\mu}=0.$ (5.15)
So, one has to study the local $s$-cohomology of ghost number zero on the jet space of $\overline{h}_{\mu\nu}$ and its derivatives. This is done in Appendix B. The result is that the $s$-cohomology is generated by the derivatives of the antisymmetrised combination $\partial_{[\mu}\,\overline{h}_{\varrho]\nu}$, which is the linearised torsion $\mathring{T}_{\mu\nu\varrho}$. Thus, one has to consider the most general action quadratic in $\mathring{T}_{\mu\nu\varrho}$, which is a combination of the three independent quadratic scalar contractions $\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\nu\varrho}$, $\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\varrho\nu}$, $\mathring{T}^{\mu\nu}{}_{\nu}\,\mathring{T}_{\mu\varrho}{}^{\varrho}$ with arbitrary coefficients, and this is the action (5.13).
The case of General Relativity (4.2) can be obtained directly by varying the massless Fierz-Pauli action (2.11), which is the most general quadratic action in the derivatives of the symmetric part $h_{\mu\nu}$, invariant under linearised diffeomorphisms
$s\,h_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}+\partial_{\nu}\,\mathring{\xi}_{\mu},\quad s\,\mathring{\xi}^{\mu}=0.$ (5.16)
Indeed, the symmetric part is the only propagating part of $\overline{h}_{\mu\nu}$ in General Relativity. If one starts with $\overline{h}_{\mu\nu}$, the antisymmetric $b_{\mu\nu}$ can be eliminated by modifying the brst rules (5.15). One introduces a new brst differential operator $s_{\textsc{fp}}$, adding an antisymmetric shift in the $b_{\mu\nu}$ transformation
$s_{\textsc{fp}}\,h_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}+\partial_{\nu}\,\mathring{\xi}_{\nu},\quad s_{\textsc{fp}}\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\nu}+2\,\mathring{\Omega}_{\nu\mu},\quad s_{\textsc{fp}}\,\mathring{\xi}_{\mu}=0,\quad s_{\textsc{fp}}\,\mathring{\Omega}_{\mu\nu}=0,$ (5.17)
or equivalently, putting the first two transformations together,
$s_{\textsc{fp}}\,\overline{h}_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}+\mathring{\Omega}_{\nu\mu},\quad s_{\textsc{fp}}\,\mathring{\xi}_{\mu}=0,\quad s_{\textsc{fp}}\,\mathring{\Omega}_{\mu\nu}=0,$ (5.18)
where $\mathring{\Omega}_{\mu\nu}=-\mathring{\Omega}_{\nu\mu}$ is anticommuting. Redefining $\tilde{\Omega}_{\nu\mu}=:\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu}+2\,\mathring{\Omega}_{\nu\mu}$, one realises that the transformation of the antisymmetric part becomes $s_{\textsc{fp}}\,b_{\mu\nu}=\tilde{\Omega}_{\mu\nu}$, $s_{\textsc{fp}}\,\tilde{\Omega}_{\mu\nu}=0$, so that $(b_{\mu\nu},\tilde{\Omega}_{\nu\mu})$ is a trivial doublet, and the local $s_{\textsc{fp}}$-cohomology does not depend on it Brandt:1989rd ; Piguet:1995er . On the other hand, the transformation of the symmetric part is the same as before $s_{\textsc{fp}}\,h_{\mu\nu}=s\,h_{\mu\nu}$. Therefore, the local $s_{\textsc{fp}}$-cohomology on the jet space of $\{h_{\mu\nu},\mathring{\xi}_{\mu},b_{\mu\nu},\tilde{\Omega}_{\mu\nu}\}$ and their derivatives is equivalent to the local $s$-cohomology on the jet space of the starting field space $\{h_{\mu\nu},\mathring{\xi}_{\mu}\}$.
The linearised torsion $\mathring{T}_{\mu\nu\varrho}$ is not invariant under the modified transformations, but it transforms with respect to $\mathring{\Omega}_{\mu\nu}$ as a gauge two-form with a spectator index
$s_{\textsc{fp}}\,\mathring{T}_{\mu\nu\varrho}=\partial_{\mu}\,\mathring{\Omega}_{\nu\varrho}-\partial_{\varrho}\,\mathring{\Omega}_{\nu\mu}.$ (5.19)
Requiring the arbitrary combination in (5.13) to be invariant with respect to the above transformations, one selects the peculiar combination (4.2), which indeed gives the Einstein-Hilbert action in the non-linear case. Consistently, the action (5.13) with coefficients (4.2) does not depend on the antisymmetric part $b_{\mu\nu}$ by integrating by parts.
6 Embedding covariant fractons in linearised MHS theory
In this section we show that all the solutions of the equations of motion of covariant fracton gauge theory (2.5) or (2.9) are solutions of the equations of motion of linearised mhs theory (5.5)–(5.6).
The space of the solutions of covariant fracton gauge theory modulo the gauge invariance (fracton moduli space) is the submanifold $\mathscr{M}_{\text{fr}}$ in the space of fields identified by the equations of motion (2.5) or (2.9),
$\mathscr{M}_{\text{fr}}=\{f^{\text{(fr)}}_{\mu\nu\varrho}(x)\,:\,E^{\text{(fr)}}_{\mu\nu}(f^{\text{(fr)}})=0\}$ (6.1)
where the elements in $\mathscr{M}_{\text{fr}}$ are identified by the invariant fracton field strength because of the gauge invariance of the equations of motion: two solutions $h^{\text{(fr)}}_{\mu\nu}(x)$ and $\tilde{h}^{\text{(fr)}}_{\mu\nu}(x)$ which differ by a fracton gauge transformation are identified in $\mathscr{M}_{\text{fr}}$. Similarly, the moduli space of the linearised mhs theory is the submanifold $\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}$ of the field space identified by the equations of motion (5.5)–(5.6):
$\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}=\{\mathring{T}_{\mu\nu\varrho}(x)\,:\,\mathring{E}^{\text{(sym)}}_{\mu\nu}(\mathring{T})=0,\;\mathring{E}^{\text{(anti)}}_{\mu\nu}(\mathring{T})=0\}.$ (6.2)
Consider an arbitrary fixed element $f_{\mu\nu\varrho}^{\text{(fr)}}$ in $\mathscr{M}_{\text{lin}}$. If we consider a linearised torsion $\tilde{T}_{\mu\nu\varrho}$, such that
$\tilde{T}_{\mu\nu\varrho}=\tfrac{1}{2}\,f_{\mu\varrho\nu}^{\text{(fr)}},$ (6.3)
which means, in particular, that we are considering a linearised torsion with vanishing totally antisymmetric part, then $\tilde{T}_{\mu\nu\varrho}$ is an element in $\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}$. To see why, consider that
$\ignorespaces\mathring{E}_{\mu\nu}^{\text{(sym)}}(\tilde{T})$ $=\alpha_{3}\,(R^{(\text{fr})}_{\mu\nu}-\tfrac{1}{2}\,R^{(\text{fr})}\,\eta_{\mu\nu})+\tfrac{1}{4}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\partial_{\alpha}\,f^{(\text{fr})\,\alpha}{}_{(\mu\nu)},$ (6.4a)
$\mathring{E}_{\mu\nu}^{\text{(anti)}}(\tilde{T})$ $=\tfrac{1}{4}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\partial_{\alpha}\,f^{(\text{fr})\,\alpha}{}_{[\mu\nu]},$ (6.4b)
where $R^{(\text{fr})}_{\mu\nu}$ and $R^{(\text{fr})}$ are given by (2.7) and (2.8). Notice that the constants $\alpha_{1}$ and $\alpha_{2}$ appear only in the combination $2\,\alpha_{1}+\alpha_{2}$, which means that, had we varied the action with, say, $\alpha_{2}=0$, we would have obtained the same result. This is not surprising, since there are only two independent quadratic scalar contractions if the totally antisymmetric part of the torsion is vanishing.
The projected symmetric equation (6.4a) corresponds to the equations of motion in covariant fracton gauge theory (2.9)
$\mathring{E}_{\mu\nu}^{\text{(sym)}}(\tilde{T})=E_{\mu\nu}^{\text{(fr)}}(f^{\text{(fr)}})=0,$ (6.5)
upon renaming the constant coefficients
$\alpha=\tfrac{1}{4}\,(2\,\alpha_{1}+\alpha_{2}),\quad\beta=\tfrac{\alpha_{3}}{2}.$ (6.6)
Nevertheless, in linearised mhs theory there is also the variation with respect to the antisymmetric part of the linearised vielbein, which is responsible for the antisymmetric part of the equations of motion (6.4b). But, excluding the case (4.2), which brings to linearised General Relativity and which is such that $02\,\alpha_{1}+\alpha_{2}+\alpha_{3}=$ 0, (6.4b) is implied by (6.4a). To show why, consider the first divergence the fracton equations of motion (2.12). If we define the following one-form Blasi:2022mbl
$j=j_{\mu}\,\mathrm{d}x^{\mu}=\partial^{\nu}\,h^{\text{(fr)}}_{\nu\mu}\,\mathrm{d}x^{\mu},$ (6.7)
the equation (2.12) can be written in this form
$\partial^{\mu}\,E_{\mu\nu}^{\text{(fr)}}\,\mathrm{d}x^{\nu}=-\mathrm{d}^{\dagger}\,\mathrm{d}\,j=0,$ (6.8)
where $\mathrm{d}$ and $\mathrm{d}^{\dagger}$ are the De Rham differential and codifferential, as it follows by explicit evaluation
$\partial^{\mu}\,\partial^{\varrho}\,f^{\text{(fr)}}_{\varrho(\mu\nu)}\,\mathrm{d}x^{\nu}$ $=2\,\partial^{\mu}\,\partial^{\varrho}\,\partial_{\varrho}\,h^{\text{(fr)}}_{\mu\nu}\,\mathrm{d}x^{\nu}-\partial^{\mu}\,\partial^{\varrho}\,\partial_{\mu}\,h^{\text{(fr)}}_{\varrho\nu}\,\mathrm{d}x^{\nu}-\partial^{\mu}\,\partial^{\varrho}\,\partial_{\nu}\,h^{\text{(fr)}}_{\varrho\mu}\,\mathrm{d}x^{\nu}=$
$=\,\partial^{2}\,\partial^{\mu}\,h^{\text{(fr)}}_{\mu\nu}\,\mathrm{d}x^{\nu}-\mathrm{d}x^{\nu}\,\partial_{\nu}\,\partial^{\mu}\,\partial^{\varrho}\,h^{\text{(fr)}}_{\mu\varrho}=$
$=\partial^{2}\,j-\mathrm{d}\,\partial_{\mu}\,j^{\mu}=-(\mathrm{d}^{\dagger}\,\mathrm{d}+\mathrm{d}\,\mathrm{d}^{\dagger})\,j+\mathrm{d}\,\mathrm{d}^{\dagger}\,j=-\mathrm{d}^{\dagger}\,\mathrm{d}\,j.$ (6.9)
Now, the equation (6.8) is equivalent to say that $j$ is $\mathrm{d}$-closed,
$\partial^{\mu}\,E_{\mu\nu}^{\text{(fr)}}\,\mathrm{d}x^{\nu}=-\mathrm{d}^{\dagger}\,\mathrm{d}\,j=0\Leftrightarrow\mathrm{d}\,j=0,$ (6.10)
[…]
$\mathring{E}_{\mu\nu}^{\text{(anti)}}(\tilde{T})\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}=-\tfrac{1}{2}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\mathrm{d}\,j,$ (6.11)
since
$\partial^{\varrho}\,f^{\text{(fr)}}_{\varrho\mu\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ $=\partial^{\varrho}\,\partial_{\varrho}\,h^{\text{(fr)}}_{\mu\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}-\partial^{\varrho}\,\partial_{\mu}\,h^{\text{(fr)}}_{\varrho\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}=-\mathrm{d}\,j.$ (6.12)
In conclusion, we have shown that
$\mathring{E}_{\mu\nu}^{\text{(anti)}}(\tilde{T})=0\Leftrightarrow\partial^{\mu}\,E_{\mu\nu}^{\text{(fr)}}(f^{(\text{fr})})=0,$ (6.13)
which means that the projected symmetric equation (6.4a) implies the antisymmetric one (6.4b). As a consequence, all the solutions of the covariant fracton gauge theory are solutions of the linearised mhs theory:
$\mathscr{M}_{\text{fr}}\subset\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}.$ (6.14)
7 BRST formulation
In this section we formulate the brst rules (2.1) defining the covariant fracton gauge theory in an equivalent way, showing in which sense the covariant fracton gauge theory can be deduced from linearised mhs theory.
Introduce two doublets $(B_{\mu\nu},\tilde{b}_{\mu\nu})$ and $(r_{\mu},\tilde{\xi}_{\mu})$, with the following transformations
$s\,B_{\mu\nu}=\tilde{b}_{\mu\nu},\quad s\,\tilde{b}_{\mu\nu}=0,$ (7.1a)
$s\,r_{\mu}=\tilde{\xi}_{\mu},\quad s\,\tilde{\xi}_{\mu}=0,$ (7.1b)
requiring that both $\tilde{b}_{\mu\nu}$ and $r_{\mu}$ are commuting with ghost number zero, so that $B_{\mu\nu}$ is anticommuting with ghost number $-1$ and $\tilde{\xi}_{\mu}$ is anticommuting with ghost number one. The local $s$-cohomology on the jet space $\{h^{(\text{fr})}_{\mu\nu},\lambda\}$ is equivalent to the local $s$-cohomology on the enlarged space including the two doublets $\{h^{(\text{fr})}_{\mu\nu},\lambda,B_{\mu\nu},\tilde{b}_{\mu\nu},r_{\mu},\tilde{\xi}_{\mu}\}$. If we redefine $\tilde{b}_{\mu\nu}$ and $\tilde{\xi}_{\mu}$ in this way
$\tilde{b}_{\mu\nu}=:b_{\mu\nu}-\partial_{\mu}\,r_{\nu}+\partial_{\nu}\,r_{\mu},\quad\tilde{\xi}_{\mu}=:\mathring{\xi}_{\mu}-\tfrac{1}{2}\,\partial_{\mu}\,\lambda,$ (7.2)
for some $b_{\mu\nu}$ and $\mathring{\xi}_{\mu}$, the last one being anticommuting, then the consistency of the brst transformations implies that
$s\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},\quad s\,\mathring{\xi}_{\mu}=0,$ (7.3)
Notice that
$s\,(h^{(\text{fr})}_{\mu\nu}+\partial_{\mu}\,r_{\nu}+\partial_{\nu}\,r_{\mu})=\partial_{\mu}\,\mathring{\xi}_{\nu}+\partial_{\nu}\,\mathring{\xi}_{\mu}$ (7.4)
is the same transformation of the linearised metric $h_{\mu\nu}$ under linearised diffeomorphisms, with ghost $\mathring{\xi}_{\mu}$. So, upon defining
$h_{\mu\nu}:=h_{\mu\nu}^{\text{(fr)}}+\partial_{\mu}\,r_{\nu}+\partial_{\nu}\,r_{\mu},$ (7.5)
the brst rules (2.1) are equivalent to the transformations of linearised mhs theory (5.15), together with the transformation of a one-form field $r_{\mu}$, whose ghost is the fracton ghost $\lambda$, shifted by the diffeomorphism ghost $\mathring{\xi}_{\mu}$, and with the transformation of a two-form field $B_{\mu\nu}$, whose transformation is the antisymmetrised derivatives of $r_{\mu}$, shifted by $b_{\mu\nu}$:
$\begin{cases}s\,\lambda=0,\\ s\,h_{\mu\nu}^{(\text{fr})}=\partial_{\mu}\,\partial_{\nu}\,\lambda\end{cases}\Leftrightarrow\begin{cases}s\,\mathring{\xi}_{\mu}=0,\quad s\,\lambda=0,\\ s\,h_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}+\partial_{\nu}\,\mathring{\xi}_{\mu},\\ s\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},\\ s\,r_{\mu}=\mathring{\xi}_{\mu}-\tfrac{1}{2}\,\partial_{\mu}\,\lambda,\\ s\,B_{\mu\nu}=b_{\mu\nu}-\partial_{\mu}\,r_{\nu}+\partial_{\nu}\,r_{\mu}.\end{cases}$ (7.6)
The local $s$-cohomology on $\{h_{\mu\nu}^{\text{(fr)}},\lambda\}$ is equivalent to that on $\{B_{\mu\nu},h_{\mu\nu},b_{\mu\nu},r_{\mu},\mathring{\xi}_{\mu},\lambda\}$. In other words, the doublet $(r_{\mu},\tilde{\xi}_{\mu})$ implements the condition “$\mathring{\xi}_{\mu}=\tfrac{1}{2}\,\partial_{\mu}\,\lambda$”, which is implied by “$r_{\mu}=0$”; and the doublet $(B_{\mu\nu},\tilde{b}_{\mu\nu})$ implements the condition “$b_{\mu\nu}=\partial_{\mu}\,r_{\nu}-\partial_{\nu}\,r_{\mu}$”, which is implied by “$B_{\mu\nu}=0$”. In this sense, the covariant fracton gauge transformations can be thought as the restriction of linearised diffeomorphisms to the longitudinal ones, as they were firstly introduced Blasi:2022mbl , and covariant fracton gauge theory can be thought as the restriction of the linearised mhs theory in the case in which the antisymmetric part of the linearised vielbein is $\mathrm{d}$-exact. As a consequence, the linearised torsion is cohomologically equivalent to the fracton field strength in this formulation. Indeed, the linearised vielbein reads
$\overline{h}_{\mu\nu}:=\tfrac{1}{2}\,(h_{\mu\nu}+b_{\mu\nu})=\tfrac{1}{2}\,h_{\mu\nu}^{\text{(fr)}}+\partial_{\mu}\,r_{\nu}+\tfrac{1}{2}\,s\,B_{\mu\nu},$ (7.7)
so that the linearised torsion becomes
$\mathring{T}_{\mu\varrho\nu}=\partial_{[\mu}\,\overline{h}_{\nu]\varrho}=\tfrac{1}{2}\,f_{\mu\nu\varrho}^{\text{(fr)}}+s\,(\tfrac{1}{2}\,\partial_{[\mu}\,B_{\nu]\varrho}),$ (7.8)
which, in particular, implies that the totally antisymmetric part of the torsion is cohomologically trivial
$\mathring{t}_{\mu\nu\varrho}=s\,(\tfrac{1}{6}\,\partial_{[\mu}\,B_{\varrho\nu]}).$ (7.9)
This is the reason why the constants $\alpha_{1}$ and $\alpha_{2}$ appear only in the combination $2\,\alpha_{1}+\alpha_{2}$ n embedding covariant fracton gauge theory in linearised mhs theory.
Acknowledgments
The author is deeply indebted with F. Fecit and C. Imbimbo for encouragements and illuminating discussions and observations.
Appendix A Weitzenböck geometry within Cartan formalism
In the formalism introduced by Cartan, one considers the vielbein one-form $e^{a}=e_{\mu}{}^{a}\,\mathrm{d}x^{\mu}$ and the spin connection one-form $\omega^{a}{}_{b}=\omega_{\mu}{}^{a}{}_{b}\,\mathrm{d}x^{\mu}$, and define the torsion two-form $T^{a}=\frac{1}{2}\,T_{\mu}{}^{a}{}_{\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ and the curvature two-form $R^{a}{}_{b}=\frac{1}{2}\,R_{\mu\nu}{}^{a}{}_{b}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ in the following way:
$\ignorespaces T^{a}(e,\omega)$ $=\mathrm{d}\,e^{a}+\omega^{a}{}_{b}\,e^{b},$ (A.1a)
$R^{a}{}_{b}(\omega)$ $=\mathrm{d}\,\omega^{a}{}_{b}+\omega^{a}{}_{c}\,\omega^{c}{}_{b}.$ (A.1b)
$\omega_{\mu}{}^{a}{}_{b}$ serves as a frame connection (for flat indices). The basis or affine connection $\Gamma_{\mu}{}^{\varrho}{}_{\nu}$ (for curved indices) is related to $\omega_{\mu}{}^{a}{}_{b}$ by the requirement that the full covariant derivative of the vielbein, both with respect to $\omega_{\mu}{}^{a}{}_{b}$ and with respect to $\Gamma_{\mu}{}^{\varrho}{}_{\nu}$ (“second vielbein postulate”), vanishes [^9] This is equivalent to define the covariant derivative with respect to the affine connection $\nabla_{\mu}(\Gamma)$ in terms of the covariant derivative with respect to the spin connection $\nabla_{\mu}(\omega)$ in this way: $\nabla_{\mu}(\Gamma)\,(e^{\varrho}{}_{a}\,V^{a})=e^{\varrho}{}_{a}\,\nabla_{\mu}(\omega)\,V^{a}$, for any $V^{a}$. Indeed on a side $\nabla_{\mu}(\Gamma)\,(e^{\varrho}{}_{a}\,V^{a})=\partial_{\mu}\,V^{\varrho}+\Gamma_{\mu}{}^{\varrho}{}_{\nu}\,V^{\nu}$; on the other $e^{\varrho}{}_{a}\,\nabla_{\mu}(\omega)\,V^{a}=\partial_{\mu}\,V^{\varrho}+e^{\varrho}{}_{a}\,(\partial_{\mu}\,e_{\nu}{}^{a}+\omega_{\mu}{}^{a}{}_{b}\,e_{\nu}{}^{b})\,V^{\nu}$, so that we obtain (A.2). VanNieuwenhuizen:1981ae :
$\partial_{\mu}\,e_{\nu}{}^{a}+\omega_{\mu}{}^{a}{}_{b}\,e_{\nu}{}^{b}-\Gamma_{\mu}{}^{\varrho}{}_{\nu}\,e_{\varrho}{}^{a}=0,$ (A.2)
which is solved by
$\Gamma_{\mu}{}^{\varrho}{}_{\nu}(e,\omega)=e^{\varrho}{}_{a}\,\partial_{\mu}\,e_{\nu}{}^{a}+\omega_{\mu}{}^{a}{}_{b}\,e^{\varrho}{}_{a}\,\,e_{\nu}{}^{b},$ (A.3)
if the vielbein is invertible. If we define
$\ignorespaces T_{\mu}{}^{\varrho}{}_{\nu}(\Gamma)$ $=\Gamma_{[\mu}{}^{\varrho}{}_{\nu]},$ (A.4a)
$R^{\varrho}{}_{\sigma\mu\nu}(\Gamma)$ $=\partial_{[\mu}\,\Gamma_{\nu]}{}^{\varrho}{}_{\sigma}+\Gamma_{[\mu}{}^{\varrho}{}_{\alpha}\,\Gamma_{\nu]}{}^{\alpha}{}_{\sigma},$ (A.4b)
then they are related to the Cartan definitions (A.1a)–(A.1b) according to
$\ignorespaces T_{\mu}{}^{\varrho}{}_{\nu}(\Gamma(e,\omega))$ $=T_{\mu}{}^{a}{}_{\nu}(e,\omega)\,e^{\varrho}{}_{a},$ (A.5a)
$R^{\varrho}{}_{\sigma\mu\nu}(\Gamma(e,\omega))$ $=R_{\mu\nu}{}^{a}{}_{b}(\omega)\,e^{\varrho}{}_{a}\,e_{\sigma}{}^{b}.$ (A.5b)
This follows straightforwardly by replacing (A.3) in (A.4a)–(A.4b).
General Relativity in usual formulation is recovered by choosing a torsionless spin connection. The spin connection appears in (A.1a) algebraically, so that the torsionless condition can be solved uniquely for the spin connection in terms of the vielbein, assuming it to be invertible. Such a spin connection, which we denote with ${\omega^{(\textsc{gr})}}_{\mu}{}^{a}{}_{b}$, reads
${\omega^{(\textsc{gr})}}_{\mu}{}^{a}{}_{b}=\tfrac{1}{2}\,e^{\nu a}\,\partial_{[\mu}\,e_{\nu]b}-\tfrac{1}{2}\,e^{\nu}{}_{b}\,\partial_{[\mu}\,e_{\nu]}{}^{a}-\tfrac{1}{2}\,e^{\nu a}\,e^{\varrho}{}_{b}\,e_{\mu}{}^{c}\,\partial_{[\nu}\,e_{\varrho]c}.$ (A.6)
The corresponding affine connection, according to (A.3), which we denote by ${\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}$, is the Levi-Civita connection:
${\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}=\Gamma_{\mu}{}^{\varrho}{}_{\nu}(e,\omega^{(\textsc{gr})})=\tfrac{1}{2}\,g^{\varrho\alpha}\,(\partial_{\mu}\,g_{\alpha\nu}+\partial_{\nu}\,g_{\mu\alpha}-\partial_{\alpha}\,g_{\mu\nu}).$ (A.7)
Indeed, using $g_{\mu\nu}=e_{\mu}{}^{a}\,e_{\nu a}$ and (A.6),
${\omega^{(\textsc{gr})}}_{\mu}{}^{a}{}_{b}\,e^{\alpha}{}_{a}\,\,e_{\beta}{}^{b}=\tfrac{1}{2}\,g^{\alpha\nu}\,(\partial_{\mu}\,g_{\nu\beta}+\partial_{\beta}\,g_{\mu\nu}-\partial_{\nu}\,g_{\mu\beta})-e^{\alpha}{}_{a}\,\partial_{\mu}\,e_{\beta}{}^{a},$ (A.8)
whence the result (A.7) follows, by using (A.3).
One could also look for a spin connection $\omega^{(\text{flat})}$ whose curvature is vanishing, instead of a torsionless spin connection. The flatness condition is uniquely solved by the Cartan-Maurer form out of an arbitrary Lorentz matrix $\lambda^{a}{}_{b}$:
${\omega^{(\text{flat})}}^{a}{}_{b}=(\lambda^{-1}\,\mathrm{d}\,\lambda)^{a}{}_{b}=\lambda_{c}{}^{a}\,\mathrm{d}\lambda^{c}{}_{b},$ (A.9)
The corresponding affine connection, according to (A.3), is again a Cartan-Maurer form: [^10] The author thanks C. Imbimbo for pointing this out.
$\Gamma_{\mu}{}^{\varrho}{}_{\nu}(e,\omega^{\text{(flat)}})=e^{\varrho}{}_{a}\,\partial_{\mu}\,e_{\nu}{}^{a}+{\omega^{(\text{flat})}}_{\mu}{}^{a}{}_{b}\,e^{\varrho}{}_{a}\,e_{\nu}{}^{b}=(\lambda_{c}{}^{b}\,e^{\varrho}{}_{b})\,\partial_{\mu}\,(\lambda^{c}{}_{a}\,e_{\nu}{}^{a})$ (A.10)
reflecting the fact that it is flat: $R(\Gamma(e,\omega^{\text{(flat)}})=0$. So, starting from a globally defined frame $e^{\mu}{}_{a}$ in a parallelisable manifold, it is always possible to make a choice of $\lambda$ such that the affine connection above is the Weitzenböck connection
$W_{\mu}{}^{\varrho}{}_{\nu}=\Gamma_{\mu}{}^{\varrho}{}_{\nu}(e,\omega^{\textsc{(w)}})=e^{\varrho}{}_{a}\,\partial_{\mu}\,e_{\nu}{}^{a}.$ (A.11)
Instead, if we start from a frame and we identity $\Gamma(e,\omega^{\text{(flat)}})$ with the Weitzenböck connection out of that frame, it means that the corresponding spin connection $\omega^{\textsc{(w)}}$ is vanishing, fixing the $\lambda$ matrix to be constant:
$\omega^{(\textsc{w})}=0.$ (A.12)
This is called Weitzenböck condition, which breaks the local Lorentz invariance, but it preserves the global Lorentz invariance.
A useful tensor is the contorsion, which is defined as the difference between an arbitrary affine connection and the Levi-Civita connection [^11] Using (A.3), we can also write the contorsion in terms of an arbitrary spin connection and the torsionless one: $-K_{\mu}{}^{\varrho}{}_{\nu}=(\omega_{\mu}{}^{a}{}_{b}-{\omega^{\textsc{(gr)}}}_{\mu}{}^{a}{}_{b})\,e^{\varrho}{}_{a}\,e_{\nu}{}^{b}$.
$-K_{\mu}{}^{\varrho}{}_{\nu}(\Gamma):=\Gamma_{\mu}{}^{\varrho}{}_{\nu}-{\Gamma^{(\textsc{gr})}}_{\mu}{}^{\varrho}{}_{\nu},$ (A.13)
which implies that $K_{\mu\varrho\nu}=K_{\mu}{}^{\sigma}{}_{\nu}\,g_{\varrho\sigma}$ is antisymmetric in the last two indices. Notice that the antisymmetric part in the first and in the last index is the torsion (A.5a):
$-K_{[\mu}{}^{\varrho}{}_{\nu]}(\Gamma)=T_{\mu}{}^{\varrho}{}_{\nu}(\Gamma).$ (A.14)
This expression can be solved for writing the contorsion in terms of the torsion
$K_{\mu\varrho\sigma}$ $=\tfrac{1}{2}\,(K_{\mu\varrho\nu}-K_{\nu\varrho\mu})+\tfrac{1}{2}\,(K_{\varrho\mu\nu}-K_{\nu\mu\varrho})+\tfrac{1}{2}\,(K_{\varrho\nu\mu}-K_{\mu\nu\varrho})=$
$=-\tfrac{1}{2}\,(T_{\mu\varrho\nu}+T_{\varrho\mu\nu}+T_{\varrho\nu\mu})=-\tfrac{1}{2}\,(T_{\mu[\varrho\nu]}+T_{\varrho\mu\nu}).$ (A.15)
We can relate the Levi-Civita connection, the Weitzenböck connection and the Weitzenböck torsion in the following way. On a side, using (A.13),
$-K_{\mu}{}^{\varrho}{}_{\nu}(W)=W_{\mu}{}^{\varrho}{}_{\nu}-{\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}.$ (A.16)
On the other, using (A),
$-K_{\mu}{}^{\varrho}{}_{\nu}(W)=\tfrac{1}{2}\,(T_{\mu}{}^{\varrho}{}_{\nu}(W)-T_{\mu\nu}{}^{\varrho}(W)+T^{\varrho}{}_{\mu\nu}(W)).$ (A.17)
Therefore,
${\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}=W_{\mu}{}^{\varrho}{}_{\nu}+\tfrac{1}{2}\,(T_{\mu}{}^{\varrho}{}_{\nu}(W)-T_{\mu\nu}{}^{\varrho}(W)+T^{\varrho}{}_{\mu\nu}(W)).$ (A.18)
Appendix B Computation of a BRST cohomology
Consider the following brst transformations:
$s\,\overline{h}_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu},\quad s\,\mathring{\xi}_{\mu}=0.$ (B.1)
$(\overline{h}_{\mu\nu},\partial_{\mu}\,\mathring{\xi}_{\nu})$ is a doublet, so it does not belong to the brst cohomology. Since
$\tfrac{1}{2}\,s\,\partial_{(\varrho}\,\overline{h}_{\mu)\nu}=\partial_{\varrho}\,\partial_{\mu}\,\mathring{\xi}_{\nu},$ (B.2a)
$s\,\mathring{T}_{\varrho\nu\mu}=0,$ (B.2b)
where $\mathring{T}_{\varrho\nu\mu}=\partial_{[\varrho}\,\overline{h}_{\mu]\nu}$, then $(\tfrac{1}{2}\,\partial_{(\varrho}\,\overline{h}_{\mu)\nu},\partial_{\varrho}\,\partial_{\mu}\,\mathring{\xi}_{\nu})$ is a doublet, whereas $\mathring{T}_{\varrho\nu\mu}$ is in the cohomology. In general,
$\tfrac{1}{(n+2)!}\,s\,\partial_{(\alpha_{n}}\,\dots\partial_{\alpha_{1}}\,\partial_{\varrho}\,\overline{h}_{\mu)\nu}=\partial_{\alpha_{n}}\dots\partial_{\alpha_{1}}\,\partial_{\varrho}\,\partial_{\mu}\,\mathring{\xi}_{\nu},$ (B.3a)
$s\,\partial_{\alpha_{n}}\,\dots\partial_{\alpha_{1}}\,\mathring{T}_{\varrho\nu\mu}=0$ (B.3b)
show that $(\tfrac{1}{(n+2)!}\,\partial_{(\alpha_{n}}\,\dots\partial_{\alpha_{1}}\,\partial_{\varrho}\,\overline{h}_{\mu)\nu},\partial_{\alpha_{n}}\dots\partial_{\alpha_{1}}\,\partial_{\varrho}\,\partial_{\mu}\,\mathring{\xi}_{\nu})$ is a doublet, for each $n$. This suggests that the cohomology is generated by $\mathring{\xi}^{\mu}$, $\mathring{T}_{\varrho\nu\mu}$ and by the derivatives of the last one:
$H(s)=\{\mathring{\xi}^{\mu},\mathring{T}_{\varrho\nu\mu},\partial_{\alpha_{1}}\,\mathring{T}_{\varrho\nu\mu},\dots,\partial_{\alpha_{n}}\dots\partial_{\alpha_{1}}\,\mathring{T}_{\varrho\nu\mu},\dots\},$ (B.4)
In order to show this, consider the decomposition $\partial_{\varrho}\,\overline{h}_{\mu\nu}$ in irreducible components.
$(1)\otimes[(1)\oplus(2)_{t}\oplus(1,1)]=(1)\oplus[(1)\oplus(2,1)_{t}\oplus(3)_{t}]\oplus[(1)\oplus(2,1)_{t}\oplus(1,1,1)],$ (B.5)
which is realised by
$\partial_{\varrho}\,\overline{h}_{\mu\nu}=\tfrac{1}{6}\,(F_{\varrho\mu\nu}+F_{\varrho\nu\mu}+G_{\varrho\mu\nu}+B_{\nu\mu\varrho}+H_{\varrho\mu\nu}),$ (B.6)
where
$F\sim(1)\oplus(2,1)_{t},\quad G\sim(3)_{t},\quad B\sim(1)\oplus(2,1)_{t},\quad H\sim(1,1,1),$ (B.7)
and we choose $F$ and $B$ antisymmetric in the first two indices. Defining the symmetric and the antisymmetric part of $\overline{h}_{\mu\nu}$,
$\overline{h}_{\mu\nu}=\tfrac{1}{2}\,(h_{\mu\nu}+b_{\mu\nu}),\quad\text{with}\quad h_{\mu\nu}=\overline{h}_{\mu\nu}+\overline{h}_{\nu\mu},\quad b_{\mu\nu}=\overline{h}_{\mu\nu}-\overline{h}_{\nu\mu},$ (B.8)
the previous tensor can are realised by
$\ignorespaces F_{\mu\nu\varrho}$ $=\partial_{\mu}\,h_{\nu\varrho}-\partial_{\nu}\,h_{\mu\varrho},$ (B.9a)
$G_{\mu\nu\varrho}$ $=\partial_{\mu}\,h_{\nu\varrho}+\partial_{\nu}\,h_{\varrho\mu}+\partial_{\varrho}\,h_{\mu\nu}=\tfrac{1}{2}\,\partial_{(\mu}\,h_{\nu\varrho)},$ (B.9b)
$B_{\mu\nu\varrho}$ $=\partial_{\mu}\,b_{\nu\varrho}-\partial_{\nu}\,b_{\mu\varrho}-2\,\partial_{\varrho}\,b_{\mu\nu},$ (B.9c)
$H_{\mu\nu\varrho}$ $=\partial_{\mu}\,b_{\nu\varrho}+\partial_{\nu}\,b_{\varrho\mu}+\partial_{\varrho}\,b_{\mu\nu}=\tfrac{1}{2}\,\partial_{[\mu}\,b_{\nu\varrho]}.$ (B.9d)
Finally, if we decompose $\partial_{\mu}\,\mathring{\xi}_{\nu}$,
$\partial_{\mu}\,\mathring{\xi}_{\nu}=\tfrac{1}{2}\,(S_{\nu\mu}+A_{\nu\mu}),\quad\text{with}\quad S_{\nu\mu}=\partial_{\mu}\,\mathring{\xi}_{\nu}+\partial_{\nu}\,\mathring{\xi}_{\mu},\quad A_{\nu\mu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},$ (B.10)
then the brst rules of the components of $\partial_{\varrho}\,\overline{h}_{\mu\nu}$ read
$s\,F_{\mu\nu\varrho}=\partial_{\varrho}\,A_{\nu\mu},$ (B.11a)
$s\,G_{\mu\nu\varrho}=\partial_{\mu}\,S_{\nu\varrho}+\partial_{\nu}\,S_{\varrho\mu}+\partial_{\varrho}\,S_{\mu\nu},$ (B.11b)
$s\,B_{\mu\nu\varrho}=-3\,\partial_{\varrho}\,A_{\nu\mu},$ (B.11c)
$s\,H_{\mu\nu\varrho}=0.$ (B.11d)
Therefore, we recognise two independent invariants
$s\,(3\,F_{\mu\nu\varrho}+B_{\mu\nu\varrho})=0,\quad s\,H_{\mu\nu\varrho}=0,$ (B.12)
and this means that the $\mathring{T}_{\mu\nu\varrho}$ generates the cohomology, since it is a combination of the unique independent invariants:
$\mathring{T}_{\mu\nu\varrho}=\tfrac{1}{6}\,(3\,F_{\mu\varrho\nu}+B_{\mu\varrho\nu}+2\,H_{\mu\varrho\nu}).$ (B.13)
Appendix C Comparison with other conventions
In Moller1978 the following tensors are defined:
$\gamma_{\lambda\nu\mu}=-K_{\mu\lambda\nu}=e_{\lambda}{}^{a}\,\nabla_{\mu}\,e_{\nu a},\quad\varphi_{\mu}=\gamma^{\nu}{}_{\mu\nu}=-e_{\lambda}{}^{a}\,\nabla_{\mu}\,e_{\nu a}.$ (C.1)
The possible quadratic scalars contractions are chosen as
$\ignorespaces\varphi^{\mu}\,\varphi_{\mu}$ $=T^{\mu\nu}{}_{\nu}\,T_{\mu\varrho}{}^{\varrho},$ (C.2a)
$\gamma^{\mu\nu\varrho}\,\gamma_{\mu\nu\varrho}$ $=\tfrac{3}{4}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\tfrac{1}{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu},$ (C.2b)
$\gamma^{\mu\nu\varrho}\,\gamma_{\varrho\nu\mu}$ $=\tfrac{1}{4}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\tfrac{1}{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu},$ (C.2c)
so that the action
$S_{\text{M{\o}ller}}=\int\mathrm{d}^{d}x\,e\,(\beta_{1}\,\varphi^{\mu}\,\varphi_{\mu}+\beta_{2}\,\gamma^{\mu\nu\varrho}\,\gamma_{\mu\nu\varrho}+\beta_{3}\,\gamma^{\mu\nu\varrho}\,\gamma_{\varrho\nu\mu})$ (C.3)
is equivalent to (4.1), if
$\beta_{1}=\alpha_{3},\quad\beta_{2}=2\,\alpha_{1}-\alpha_{2},\quad\beta_{3}=3\,\alpha_{2}-2\,\alpha_{1},$ (C.4)
and the Einstein-Hilbert action is recovered if
$\beta_{1}=-1,\quad\beta_{2}=0,\quad\beta_{3}=1.$ (C.5)
In Hayashi:1979qx , the following tensors are defined in 4d:
$\ignorespaces{T^{\textsc{(hs)}}}^{\lambda}{}_{\mu\nu}$ $=T_{\nu}{}^{\lambda}{}_{\mu}=e^{\lambda}{}_{a}\,\partial_{[\nu}\,e_{\mu]}{}^{a},$ (C.6a)
$v_{\mu}$ $={T^{\textsc{(hs)}}}^{\lambda}{}_{\lambda\mu}=T_{\mu\lambda}{}^{\lambda},$ (C.6b)
$a_{\mu}$ $=\tfrac{1}{6}\,\varepsilon_{\mu\nu\varrho\sigma}\,{T^{\textsc{(hs)}}}^{\nu\varrho\sigma}=\tfrac{1}{6}\,\varepsilon_{\mu\nu\varrho\sigma}\,T^{\sigma\nu\varrho},$ (C.6c)
$t^{\textsc{(hs)}}_{\lambda\mu\nu}$ $=\tfrac{1}{2}\,({T^{\textsc{(hs)}}}_{\lambda\mu\nu}+{T^{\textsc{(hs)}}}_{\mu\lambda\nu})+\tfrac{1}{6}\,(g_{\nu\lambda}\,v_{\mu}+g_{\nu\mu}\,v_{\lambda})-\tfrac{1}{3}\,g_{\lambda\mu}\,v_{\nu}=$
$=\tfrac{1}{2}\,(T_{\nu\lambda\mu}+T_{\nu\mu\lambda})+\tfrac{1}{6}\,(g_{\nu\lambda}\,T_{\mu\varrho}{}^{\varrho}+g_{\nu\mu}\,T_{\lambda\varrho}{}^{\varrho})-\tfrac{1}{3}\,g_{\lambda\mu}\,T_{\nu\varrho}{}^{\varrho}.$ (C.6d)
$v_{\mu}$, $a_{\mu}$ and $t_{\lambda\mu\nu}$ are the irreducible components of the torsion in 4d, with $a_{\mu}$ (4d dual to) the totally antisymmetric part. In particular, $t_{\lambda\mu\nu}$ is traceless with vanishing totally antisymmetric part, and it is chosen to be symmetric in the first two indices. The choice for a basis for the quadratic scalar contractions of the torsion is the following:
$\ignorespaces{t^{\textsc{(hs)}}}^{\lambda\mu\nu}\,t^{\textsc{(hs)}}_{\lambda\mu\nu}$ $=\tfrac{1}{2}\,T^{\lambda\mu\nu}\,T_{\lambda\mu\nu}+\tfrac{1}{2}\,T^{\lambda\mu\nu}\,T_{\lambda\nu\mu}-\tfrac{1}{2}\,T^{\lambda\mu}{}_{\mu}\,T_{\lambda\nu}{}^{\nu},$ (C.7a)
$v^{\mu}\,v_{\mu}$ $=T^{\lambda\mu}{}_{\mu}\,T_{\lambda\nu}{}^{\nu},$ (C.7b)
$a^{\mu}\,a_{\mu}$ $=\tfrac{1}{18}\,T^{\lambda\mu\nu}T_{\lambda\mu\nu}+\tfrac{1}{9}\,T^{\lambda\mu\nu}\,T_{\lambda\nu\mu},$ (C.7c)
so that the action
$S_{\textsc{HS}}$ $=\int\mathrm{d}^{4}x\,e\,(a_{1}\,{t^{\textsc{(hs)}}}^{\lambda\mu\nu}\,t^{\textsc{(hs)}}_{\lambda\mu\nu}+a_{2}\,v^{\mu}\,v_{\mu}+a_{3}\,a^{\mu}\,a_{\mu})=$
$=\int\mathrm{d}^{4}x\,e\,[(c_{1}-\tfrac{1}{3})\,{t^{\textsc{(hs)}}}^{\lambda\mu\nu}\,t^{\textsc{(hs)}}_{\lambda\mu\nu}+(c_{2}+\tfrac{1}{3})\,v^{\mu}\,v_{\mu}+(c_{3}-\tfrac{3}{4})\,a^{\mu}\,a_{\mu}]$ (C.8)
is equivalent to (4.1) if
$\ignorespaces a_{1}$ $=c_{1}-\tfrac{1}{3}=\tfrac{2}{3}\,(2\,\alpha_{1}+\alpha_{2}),$ (C.9a)
$a_{2}$ $=c_{2}+\tfrac{1}{3}=\tfrac{1}{3}\,(2\,\alpha_{1}+\alpha_{2}+3\,\alpha_{3}),$ (C.9b)
$a_{3}$ $=c_{3}-\tfrac{3}{4}=6\,(\alpha_{1}-\alpha_{2}),$ (C.9c)
and the Hilbert-Einstein action is obtained if
$a_{1}=-\tfrac{1}{3},\quad a_{2}=\tfrac{1}{3},\quad a_{3}=-\tfrac{3}{4},$ (C.10)
or equivalently,
$c_{1}=c_{2}=c_{3}=0.$ (C.11)

### v3

@@
 Abstract
-The relation between covariant fracton gauge theory and Møller-Hayashi-Shirafuji extension of teleparallel gravity is investigated.
-The former is the gauge theory of a rank two symmetric tensor with gauge symmetry given by the double derivative of a scalar parameter; the latter is a class of theories which generalises Einstein gravity, whose action is quadratic in the Weitzenböck torsion.
-We show that the solutions of covariant fracton gauge theory can be viewed as solutions of Møller-Hayashi-Shirafuji theory, providing a new insight in the relation between covariant fractons and gravity, and elucidating the meaning of covariant fracton theories as a new type of gauge theory.
+The relation between covariant fracton gauge theory and Møller-Hayashi-Shirafuji theory of gravity is investigated.
+The former is the gauge theory of a rank-two symmetric tensor with gauge symmetry given by the double derivative of a scalar parameter; the latter is the most general theory, whose action is quadratic in the Weitzenböck torsion.
+We show that the solutions of covariant fracton gauge theory describe a subsector of the space of solutions of Møller-Hayashi-Shirafuji theory, providing a new insight in the relation between covariant fractons and gravity, and elucidating the meaning of covariant fracton theory as a new type of gauge theory.
 1 Introduction and outlook
-In this note we pursue the investigation on covariant fracton gauge theory and its properties, studying the relation between covariant fractons and a generalisation of teleparallel gravity.
-Fracton phases, which are excitations characterised by restricted mobility due to the conservation of the dipole momentum Pretko:2017xar ; Seiberg:2020bhn ; Burnell:2021reh , were firstly introduced in condensed matter physics Vijay:2015mka , but attracted interest in a broader context.
-In particular, covariant fracton gauge theory, introduced for the first time in Blasi:2022mbl , is the gauge theory of a rank two symmetric tensor with gauge symmetry given by the double derivative of a scalar parameter.
+In this paper we pursue the investigation on covariant fracton gauge theory and its properties, studying the relation between covariant fractons and teleparallel gravity.
+Fracton phases, which are excitations characterised by restricted mobility due to the conservation of the dipole moment Pretko:2017xar ; Seiberg:2020bhn ; Burnell:2021reh , were firstly introduced in condensed matter physics Vijay:2015mka , but attracted interest in a broader context.
+In particular, covariant fracton gauge theory, introduced for the first time in Blasi:2022mbl , is the gauge theory of a rank-two symmetric tensor $h_{\mu\nu}^{(\text{fr})}$ with gauge symmetry given by the double derivative of a scalar parameter
+$\delta\,h^{\text{(fr)}}_{\mu\nu}=\partial_{\mu}\,\partial_{\nu}\,\lambda.$ (1.1)
 Covariant fracton gauge theory elucidates the main properties of ordinary fracton excitations, in the context of the higher rank extensions of Maxwell theory Bertolini:2022ijb , and it can be thought as an extension of linearised Einstein gravity, invariant with respect to longitudinal diffeomorphisms Bertolini:2023juh ; Afxonidis:2023pdq .
-In particular, the stability of fracton gravity Afxonidis:2024tph , three-dimensional Chern-Simons-like models with covariant fracton gauge symmetry Bertolini:2024yur , and the analogue of the electromagnetic duality in covariant fracton gauge theory Bertolini:2025jov have been recently investigated.
-The most general quadratic action in the derivatives of the fracton gauge field $h_{\mu\nu}^{\text{(fr)}}$ can be written in terms of the two possible scalar contractions of the gauge invariant fracton field strength Rovere:2024nwc  [^1] It appears also in Bertolini:2024apg at the linear level.
-$h_{\mu\nu}^{\text{(fr)}}$ and $f_{\mu\nu\varrho}^{(\text{fr})}$ are respectively denoted with $h_{\mu\nu}$ and $f_{\mu\nu\varrho}$ in Rovere:2024nwc .
+Recently, three-dimensional Chern-Simons-like models with covariant fracton gauge symmetry Bertolini:2024yur , and the analogue of the electromagnetic duality in covariant fracton gauge theory Bertolini:2025jov have been investigated.
+[…]
 $f^{\text{(fr)}}_{\mu\nu\varrho}=\partial_{\mu}\,h_{\nu\varrho}^{\text{(fr)}}-\partial_{\nu}\,h_{\mu\varrho}^{\text{(fr)}},$ (1.2)
-which is analogous to the field strength in electromagnetism. So, the action of the theory reads
-$S_{\text{fr}}=\int\mathrm{d}^{d}x\,(\alpha\,{f^{\text{(fr)}}}^{\mu\nu\varrho}\,f^{\text{(fr)}}_{\mu\nu\varrho}+\beta\,{f^{\text{(fr)}}}^{\mu\nu}{}_{\nu}\,{f^{\text{(fr)}}}_{\mu\varrho}{}^{\varrho}),$ (1.2)
+which is analogous to the field strength in electromagnetism. The action reads
+[…]
 where $\alpha$ and $\beta$ are arbitrary constants.
-Taking $\alpha$ and $\beta$ such that $\alpha+\frac{1}{2}\,\beta=0$, and replacing the fracton gauge field with the parturbation of the metric in linearised gravity, one gets the Fierz-Pauli action (see (2.11)) Fierz:1939ix , which is invariant under the full group of linearised diffeomorphisms.
+Taking $\alpha$ and $\beta$ such that $\frac{\beta}{\alpha}=-2$, and replacing the fracton gauge field with the perturbation of the metric in linearised gravity, one gets the Fierz-Pauli action (see equation (2.12)) Fierz:1939ix , which is invariant under the full group of linearised diffeomorphisms.
 Since the whole dynamics of covariant fracton gauge theory is encoded in the fracton field strength, one might ask which its gravitational analogue is.
-We observe in this note that the covariant fracton field strength is the analogue of the linearised Weitzenböck torsion in gravity, when the latter has vanishing totally antisymmetric part.
-The first attempts of a torsion formulation of General Relativity, usually known as teleparallel gravity, date back to Einstein Einstein1928 , who unsuccessfully tried to unificate gravitation and electromagnetism in four dimensions, using the sixteen degrees of freedom of the vierbein, according to the works of Cartan Cartan1922 ; Cartan1923 ; Cartan1924 ; Cartan1925 and Weitzenböck Weitzenbock1928 .
-For the history and prehistory of the teleparallel gravity and its extension see for example Shirafuji:1995xc .
-In teleparallel gravity the dynamical variables of General Relativity are encoded in the torsion of a flat connection, known as Weitzenböck connection:
+We observe in this paper that the covariant fracton field strength is the analogue of the linearised Weitzenböck torsion in gravity, when the linearised vielbein has vanishing totally antisymmetric part.
+The first attempts of a torsion formulation of General Relativity, usually known as teleparallel equivalent of General Relativity, date back to Einstein Einstein1928 , who unsuccessfully tried to unificate gravitation and electromagnetism in four dimensions, using the sixteen degrees of freedom of the vierbein, according to Cartan’s Cartan1922 ; Cartan1923 ; Cartan1924 ; Cartan1925 and Weitzenböck’s Weitzenbock1928 works.
+For the history and prehistory of the teleparallel gravity see for example Shirafuji:1995xc ; Maluf:2013gaa .
+In teleparallel gravity the dynamical variables are encoded in the torsion of a flat connection, known as Weitzenböck connection:
 $W_{\mu}{}^{\varrho}{}_{\nu}=e^{\varrho}{}_{a}\,\partial_{\mu}\,e_{\nu}{}^{a}$. (1.4)
-The notion of Weitzenböck torsion is very useful for its geometrical implications, since it was employed in the geometrisation of gauge symmetries in the context of Scherk-Schwarz compactifications Scherk:1979zr , and in generalised geometry Coimbra:2011ky ; Aldazabal:2013mya ; Lee:2014mla .
-Møller studied extensions of teleparallel gravity, by considering the wider class of theories defined by the most general action quadratic in the Weitzenböck torsion Moller1961a ; Moller1961b ; Moller1978 .
-At the same time, independently, the possible phenomenological implications of the model were firstly studied by Hayashi and Shirafuji Hayashi:1979qx .
+where $e_{\mu}{}^{a}$ is the vielbein, and $e^{\mu}{}_{a}$ its inverse. The Weitzenböck torsion, which is the antisymmetric part of the previous connection
+$T_{\mu}{}^{\varrho}{}_{\nu}=W_{[\mu}{}^{\varrho}{}_{\nu]}=e^{\varrho}{}_{a}\,\partial_{[\mu}\,e_{\nu]}{}^{a}$, (1.5)
+is a well-defined covariant tensor, which is very useful for its geometrical implications: namely, it was employed in the geometrisation of gauge symmetries in the context of Scherk-Schwarz compactifications Scherk:1979zr , and in generalised geometry Coimbra:2011ky ; Aldazabal:2013mya ; Lee:2014mla .
+Møller studied extensions of General Relativity in teleparallel formulation, by considering the wider class of theories defined by the most general action quadratic in the Weitzenböck torsion Moller1961a ; Moller1961b ; Moller1978 .
+At the same time, independently, the possible phenomenological implications of the model were first studied by Hayashi and Shirafuji Hayashi:1979qx .
 The model can be also viewed as the gauge theory of translations Hayashi:1967se .
 The action for the theory considered by Møller and by Hayashi and Shirafuji, which we will call Møller-Hayashi-Shirafuji theory (mhs), reads
-$S_{\textsc{mhs}}=\frac{1}{2}\int\mathrm{d}^{d}x\,e\,(\alpha_{1}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\alpha_{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}+\alpha_{3}\,T^{\mu\nu}{}_{\nu}\,T_{\mu\varrho}{}^{\varrho}),$ (1.4)
-where $T_{\mu}{}^{\varrho}{}_{\nu}=W_{[\mu}{}^{\varrho}{}_{\nu]}$ is the Weitzenböck torsion, which is a well-defined tensor, and $\alpha_{1},\alpha_{2},\alpha_{3}$ are arbitrary constants.
-We will show that all the solutions of the equations of motion of covariant fracton gauge theory are solutions of the equations of motion of linearised mhs theory.
-To make the embedding more precise, we will take into account the brst formulation Becchi:1974md ; Becchi:1974xu ; Becchi:1975nq ; Tyutin:1975qk of covariant fracton gauge theory.
-In this setup, the gauge parameter is replaced by an anticommuting scalar field (fracton ghost) $\lambda$, and the gauge transformation is realised by a nilpotent differential operator $s$ on the space of the fracton gauge field and of the ghost and their derivatives, evaluated in the same spacetime point:
-$s\,\lambda=0,$ (1.5a)
-$s\,h_{\mu\nu}^{(\text{fr})}=\partial_{\mu}\,\partial_{\nu}\,\lambda.$ (1.5b)
-The local $s$-cohomology, which classifies actions and anomalies of the theory, is generated by the undifferentiated ghost and by the fracton field strength and its derivatives, as in the analogous case of electromagnetism Rovere:2024nwc .
-We will show that the brst transformations of covariant fracton gauge theory can be conveniently formulated in an equivalent way, by adding two trivial doublets.
-The resulting transformations are
-$s\,\mathring{\xi}_{\mu}=0,$ (1.6a)
-$s\,\lambda=0,$ (1.6b)
-$s\,h_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}+\partial_{\nu}\,\mathring{\xi}_{\mu},$ (1.6c)
-$s\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},$ (1.6d)
-$s\,r_{\mu}=\mathring{\xi}-\tfrac{1}{2}\,\partial_{\mu}\,\lambda,$ (1.6e)
-$s\,B_{\mu\nu}=b_{\mu\nu}-\partial_{\mu}\,r_{\nu}+\partial_{\nu}\,r_{\mu},$ (1.6f)
-where, in particular, the symmetric rank two tensor $h_{\mu\nu}$, which replaces the fracton gauge field, shifted by the symmetrised first derivative of $r_{\mu}$,
-$h_{\mu\nu}=h_{\mu\nu}^{\text{(fr)}}+\partial_{\mu}\,r_{\nu}+\partial_{\nu}\,r_{\mu},$ (1.7)
-enjoys the gauge transformation of linearised gravity, given by linearised diffeomorphisms and generated by the anticommuting ghost $\mathring{\xi}$.
-$h_{\mu\nu}$ and $b_{\mu\nu}$, the latter being an antisymmetric rank two tensor, can be respectively viewed as the symmetric and the antisymmetric part of the linearised vielbein.
-In Einstein gravity, in which only $h_{\mu\nu}$ is propagating, local Lorentz transformations are included in order to gauge away $b_{\mu\nu}$.
-In the above transformation there is no local Lorentz transformations: this is the case of Weitzenböck geometry, which is flat and it is invariant only under global Lorentz transformations Hayashi:1977jd .
-Nevertheless, the last of the above transformations requires $b_{\mu\nu}$ to be cohomologically equivalent to a $\mathrm{d}$-exact two form (i.e.
-they differ by an $s$-exact term, which vanishes in the cohomology space).
-As we will see, the consequence is that the linearised Weitzenböck torsion is cohomologically equivalent to (one half) the fracton field strength, and so the totally antisymmetric part of the linearised Weitzenböck torsion is cohomologically trivial.
-Finally, the transformation of $r_{\mu}$ serves to implement the restriction of linearised diffeomorphisms to the “longitudinal” ones, bringing to the covariant fracton gauge transformations as they were firstly introduced Blasi:2022mbl .
-In Bertolini:2024apg a possible non-linear extension of covariant fracton gauge theory is discussed, by means of the Frölicher-Nijenhuis bracket, which acts on vector-valued differential forms and it coincides with the Lie derivative between vector fields when it acts on vectors FroNij1 ; FroNij2 ; FroNij3 .
+[…]
+where $\alpha_{1},\alpha_{2},\alpha_{3}$ are arbitrary constants.
+We will show that, when $\frac{\beta}{\alpha}\neq 2$, the solutions of the equations of motion of covariant fracton gauge theory can be viewed as the subsector in the space of solutions of the equations of motion of linearised mhs theory, identified by the vanishing antisymmetric part of the linearised vielbein, denoted with $b_{\mu\nu}$,
+$b_{\mu\nu}=0.$ (1.7)
+The embedding of covariant fracton solutions in mhs theory is realised when the fracton gauge field $h_{\mu\nu}^{(\text{fr})}$ is identified with the symmetric part of the linearised vielbein, denoted with $h_{\mu\nu}$,
+$h_{\mu\nu}^{\text{(fr)}}=h_{\mu\nu},$ (1.8)
+and the constants $\alpha$ and $\beta$ in the covariant fracton action (1.3) are identified with combinations of the constant $\alpha_{1},\alpha_{2}$, and $\alpha_{3}$ in the mhs action (1.6)
+$\alpha=\tfrac{1}{4}\,(2\,\alpha_{1}+\alpha_{2}),\quad\beta=\tfrac{1}{2}\,\alpha_{3}.$ (1.9)
+To show this, focusing on four dimensions for definiteness, we will compare the particle spectrum of both theories.
+Except for some particular cases, tuned by special values of the constants $\alpha_{1},\alpha_{2}$, and $\alpha_{3}$, such that the gauge invariance is enlarged, the spectrum of mhs theory in four dimensions consists in six degrees of freedom, given by a particle with helicity $\pm 2$, one with helicity $\pm 1$, and two scalar particles.
+The antisymmetric field $b_{\mu\nu}$ describes only a scalar particle (as expected, since a two-form gauge field in four dimensions is dual to a scalar field).
+The sector with vanishing $b_{\mu\nu}$ coincides with the particle spectrum of covariant fracton gauge theory, which indeed describes the remaining five propagating degrees of freedom Afxonidis:2023pdq , except for the particular cases in which the gauge symmetry is enlarged.
+Linearised diffeomorphisms act on $b_{\mu\nu}$ as a one-form gauge transformation for a two-form field (usually known as Ramond-Kalb field Kalb:1974yc ).
+The condition (1.7) is not diffeomorphism invariant, but it leaves a residual gauge symmetry when the gauge parameter is exact, that is, when it is a derivative.
+This is the case of longitudinal diffeomorphisms, so that the residual gauge symmetry is the fracton gauge symmetry.
+In a sense, covariant fracton solutions spontaneously break the diffeomorphism symmetry of mhs theory, preserving only the fracton gauge symmetry.
+To elucidate this setting, we will take into account the brst formulation Becchi:1974md ; Becchi:1974xu ; Becchi:1975nq ; Tyutin:1975qk of covariant fracton gauge theory.
+In the brst setup, the gauge parameter is replaced by an anticommuting scalar field (fracton ghost) $\lambda$, and the gauge transformation is realised by a nilpotent differential operator $s$ on the space of the fracton gauge field and of the ghost and their derivatives, evaluated in the same spacetime point.
+The local $s$-cohomology, which classifies the consistent anomalies, is generated by the undifferentiated ghost and by the fracton field strength and its derivatives, as in the analogous case of electromagnetism Rovere:2024nwc .
+We will show that the brst transformations of covariant fracton gauge theory can be conveniently formulated in an equivalent way, by including two brst trivial doublets.
+In this formulation a symmetric rank-two tensor $h_{\mu\nu}$ and an antisymmetric one $b_{\mu\nu}$ are included.
+They can be respectively viewed as the symmetric and the antisymmetric part of the linearised vielbein, but the brst transformations do not involve any local Lorentz transformation: this is the case of Weitzenböck geometry, which is flat and it is invariant only under global Lorentz transformations Hayashi:1977jd .
+Nevertheless, on the one side, the brst transformations require the restriction of linearised diffeomorphisms to the “longitudinal” ones, bringing us to covariant fracton gauge transformations as they were first introduced Blasi:2022mbl ; on the other side, $b_{\mu\nu}$ is equivalent to zero.
+This individuates the subsector of mhs theory in which covariant fracton gauge theory lives.
+Covariant fractons were introduced in Blasi:2022mbl only at linear level.
+In Afxonidis:2024tph it is shown that the theory is afflicted by instabilities, which in principle can be cured by considering suitable interactions.
+Going towards this direction, the result in this note can be viewed as a possible way to embed covariant fracton theory in a larger theory of gravity, whose non-linear extension is automatically given by the extension of General Relativity in teleparallel formulation, studied by Møller, Hayashi, and Shirafuji.
+A possible non-linear extension of covariant fracton gauge theory is also discussed in Bertolini:2024apg , where the Frölicher-Nijenhuis bracket is employed, acting on vector-valued differential forms and coinciding with the Lie derivative between vector fields when it acts on vectors FroNij1 ; FroNij2 ; FroNij3 .
 The non-linear extension of the fracton gauge transformation is realised by using a vector parameter, which is restricted to be the derivative of a scalar parameter only at the linear level; the gauge field is replaced by a vector-valued one-form, and a constraint is added in order to make it symmetric; a possible non-linear equation of motion is postulated, without an action from which to deduce it; the linearisation of such an equation is the equation of motion of covariant fracton gauge theory (with $\beta=0$ in (1.3)).
 It would be interesting to exploit the relation between covariant fractons and Weitzenböck torsion in understanding deeper this non-linear extension.
-In particular, it could be nice to see if it is possible to relate the postulated non-linear equation of motion with the equations of mhs theory, since it seems natural to identity the vector-valued one form with the linearised vielbein.
-The rest of the paper is organised as follows: in Section 2 covariant fracton theories are formulated in a way which makes manifest the relation with torsion formulation of gravity; in Section 3, 4, and 5 torsion formulation of General Relativity, the mhs theory, and its linearisation are reviewed; Section 6 is devoted to the embedding of covariant fractons in torsion gravity; finally, in Section 7 the brst formulation is investigated.
+In particular, one could try to see if it is possible to relate the postulated non-linear equation of motion with the equations of mhs theory, since it seems natural to identity the vector-valued one form with the linearised vielbein.
+The rest of the paper is organised as follows: in Section 2 covariant fracton theory is formulated in a way which makes manifest the relation with torsion formulation of gravity; in Section 3, 4, and 5 torsion formulation of General Relativity, the mhs theory, and its linearisation are reviewed; Section 6 and 7 the particle spectrum of covariant fracton and mhs theory are computed and compared, providing the embedding of covariant fractons in mhs theory; finally, in Section 8 the equivalent brst formulation is investigated.
+In Appendix A there is a summary of Weitzenböck geometry; in Appendix B a computation for brst cohomology is shown; Appendix C is devoted to a comparison between different conventions; and in Appendix D there is a summary of polarisation matrices.
 2 Covariant fracton gauge theory in “torsion-like formulation”
-In this section covariant fracton gauge theory Blasi:2022mbl is formulated in a way in which it will be natural the embedding of in gravitational theories in torsion formulation.
-Covariant fracton gauge theory is defined by a symmetric rank two tensor $h^{(\text{fr})}_{\mu\nu}$ and by a scalar gauge parameter, promoted to an anticommuting ghost $\lambda$ in brst formalism Becchi:1974md ; Becchi:1974xu ; Becchi:1975nq ; Tyutin:1975qk .
+In this section covariant fracton gauge theory Blasi:2022mbl is formulated in a way in which its embedding in gravitational theories in torsion formulation will be natural.
+Covariant fracton gauge theory is defined by a symmetric rank-two tensor $h^{(\text{fr})}_{\mu\nu}$ and by a scalar gauge parameter, promoted to an anticommuting ghost $\lambda$ in brst formalism Becchi:1974md ; Becchi:1974xu ; Becchi:1975nq ; Tyutin:1975qk .
 The nilpotent brst transformations are
@@
 $\partial_{\sigma}\,f^{\text{(fr)}}_{\mu\nu\varrho}+\partial_{\mu}\,f^{\text{(fr)}}_{\nu\sigma\varrho}+\partial_{\nu}\,f^{\text{(fr)}}_{\sigma\mu\varrho}=0.$ (2.3c)
-The most general invariant quadratic action in the derivative of $h^{(\text{fr})}_{\mu\nu}$ can be written in terms of the fracton field strength, since the local brst cohomology of ghost number zero is generated by the fracton field strength and its derivatives Rovere:2024nwc .
-There are only two independent quadratic scalar contractions of the fracton field strength, since it transforms in the mixed Young tableau $(2,1)$.
-[^2] We denote with $(\lambda_{1},\dots,\lambda_{n})$ the Young tableau with $n$ rows and $\lambda_{i}$ boxes in the $i$-th row, $i=1,\dots,n$.
+The most general invariant quadratic action in the derivative of $h^{(\text{fr})}_{\mu\nu}$ can be written in terms of the fracton field strength.
+There are only two independent quadratic scalar contractions of the fracton field strength, since it transforms in the mixed Young tableau (2,1)(2,1).
+[^2] We denote with $(\lambda_{1},\dots,\lambda_{n})$ the Young tableau with $n$ rows and $\lambda_{i}$ boxes in the $i$th row, $i=1,\dots,n$.
+$(\lambda_{1},\dots,\lambda_{n})_{t}$ is the corresponding traceless part.
 So, the most general quadratic invariant action can be written as
-$S_{\text{fr}}$ $=\int\mathrm{d}^{d}x\,(\alpha\,{f^{\text{(fr)}}}^{\mu\nu\varrho}\,f^{\text{(fr)}}_{\mu\nu\varrho}+\beta\,{f^{\text{(fr)}}}^{\mu\nu}{}_{\nu}\,{f^{\text{(fr)}}}_{\mu\varrho}{}^{\varrho})=$
+[…]
 $=\int\mathrm{d}^{d}x\,(2\,\alpha\,\partial^{\varrho}\,h^{\text{(fr)}\,\mu\nu}\,\partial_{\varrho}\,h^{\text{(fr)}}_{\mu\nu}-2\,(\alpha-\tfrac{\beta}{2})\,\partial^{\nu}\,h^{\text{(fr)}}_{\mu\nu}\,\partial_{\varrho}\,h^{\text{(fr)}\,\mu\varrho}\,+$
@@
 for arbitrary constants $\alpha$ and $\beta$. The corresponding equations of motion require the following symmetric tensor to vanish
-$E^{\text{(fr)}}_{\mu\nu}:=-\frac{1}{2}\,\frac{\delta\,S_{\text{fr}}}{\delta\,{h^{(\text{fr})}}^{\mu\nu}}=\alpha\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho}{}_{(\mu\nu)}-\tfrac{\beta}{2}\,\partial_{(\mu}\,{f^{\text{(fr)}}}_{\nu)\varrho}{}^{\varrho}+\beta\,\eta_{\mu\nu}\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho\sigma}{}_{\sigma},$ (2.5)
+$E^{\text{(fr)}}_{\mu\nu}:=-\frac{1}{2}\,\frac{\delta\,S_{\text{fr}}}{\delta\,{h^{(\text{fr})}}^{\mu\nu}}$ $=\alpha\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho}{}_{(\mu\nu)}-\tfrac{\beta}{2}\,\partial_{(\mu}\,{f^{\text{(fr)}}}_{\nu)\varrho}{}^{\varrho}+\beta\,\eta_{\mu\nu}\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho\sigma}{}_{\sigma}$= (2.5)
+$=2\,\alpha\,\partial^{2}\,h_{\mu\nu}^{\text{(fr)}}-\tfrac{2\,\alpha-\beta}{2}\,\partial_{(\mu}\,\partial^{\varrho}\,h^{\text{(fr)}}_{\nu)\varrho}-\beta\,\partial_{\mu}\,\partial_{\nu}\,h^{\text{(fr)}\,\varrho}{}_{\varrho}\,$+
+$\quad+\beta\,\eta_{\mu\nu}\,(\partial^{2}\,h^{\text{(fr)}\,\varrho}{}_{\varrho}-\partial_{\varrho}\,\partial_{\sigma}\,h^{\text{(fr)}\,\varrho\sigma})$ (2.6)
 where the round brackets among indices denote symmetrisation without any numerical coefficient.
@@
 Writing $\partial_{(\mu}\,{f^{\text{(fr)}}}_{\nu)\varrho}{}^{\varrho}$ and $\partial_{\mu}\,{f^{\text{(fr)}}}^{\mu\nu}{}_{\nu}$ in terms of the “Ricci tensor” and of the “Ricci scalar” in the equations of motion (2.5), one recognises the analogue of the linearised Einstein tensor:
-$E^{\text{(fr)}}_{\mu\nu}=2\,\beta\,(R^{\text{(fr)}}_{\mu\nu}-\tfrac{1}{2}\,R^{\text{(fr)}}\,\eta_{\mu\nu})+(\alpha+\tfrac{\beta}{2})\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho}{}_{(\mu\nu)}.$ (2.9)
-Indeed, if one replace $h^{\text{(fr)}}_{\mu\nu}$ in the action (2) with a rank two symmetric tensor $h_{\mu\nu}$, thought as the perturbation of a metric $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$, and $\alpha$ and $\beta$ are such that
-$\alpha+\tfrac{\beta}{2}=0,$ (2.10)
-choosing the normalisation $\alpha=\frac{1}{2}$, one recovers the Fierz-Pauli action Fierz:1939ix
-$S_{\textsc{fp}}=\tfrac{1}{2}\int\mathrm{d}^{d}x\,(\partial_{\varrho}\,h_{\mu\nu}\,\partial^{\varrho}\,h^{\mu\nu}-2\,\partial^{\nu}\,h_{\mu\nu}\,\partial_{\varrho}\,h^{\mu\varrho}+2\,\partial_{\mu}\,h_{\nu}{}^{\nu}\,\partial_{\varrho}\,h^{\mu\varrho}-\partial_{\mu}\,h_{\nu}{}^{\nu}\,\partial^{\mu}\,h_{\varrho}{}^{\varrho}).$ (2.11)
+$E^{\text{(fr)}}_{\mu\nu}=2\,\beta\,(R^{\text{(fr)}}_{\mu\nu}-\tfrac{1}{2}\,R^{\text{(fr)}}\,\eta_{\mu\nu})+(\alpha+\tfrac{\beta}{2})\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho}{}_{(\mu\nu)}$.
+(2.10)
+Indeed, if one replace $h^{\text{(fr)}}_{\mu\nu}$ in the action (2) with a rank-two symmetric tensor $h_{\mu\nu}$, thought as the perturbation of a metric $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$, and $\alpha$ and $\beta$ are such that
+$\tfrac{\beta}{\alpha}=-2,$ (2.11)
+choosing the normalisation $\alpha=\frac{1}{2}$, one recovers the massless Fierz-Pauli action Fierz:1939ix
+[…]
 Consistently, imposing the condition (2.11) in the equations of motion (2.10), the last term vanishes and one gets the linearised Einstein equations.
@@
 $\ignorespaces\nabla_{\varrho}(\Gamma)\,g_{\mu\nu}$ $=\partial_{\varrho}\,g_{\mu\nu}-\Gamma_{\varrho}{}^{\sigma}{}_{\mu}\,g_{\sigma\nu}-\Gamma_{\varrho}{}^{\sigma}{}_{\nu}\,g_{\mu\sigma}=0,$ (3.1a)
-$T_{\mu}{}^{\varrho}{}_{\nu}(\Gamma)$ $=\Gamma_{\mu}{}^{\varrho}{}_{\nu}-\Gamma_{\nu}{}^{\varrho}{}_{\mu}=0,$ (3.1b)
-$R_{\mu\nu}{}^{\varrho}{}_{\sigma}(\Gamma)$ $=\partial_{\mu}\,\Gamma_{\nu}{}^{\varrho}{}_{\sigma}-\partial_{\nu}\,\Gamma_{\mu}{}^{\varrho}{}_{\sigma}+\Gamma_{\mu}{}^{\varrho}{}_{\tau}\,\Gamma_{\nu}{}^{\tau}{}_{\sigma}-\Gamma_{\nu}{}^{\varrho}{}_{\tau}\,\Gamma_{\mu}{}^{\tau}{}_{\sigma}.$ (3.1c)
+[…]
+[…]
 As known, there is a unique choice for such a connection, the Levi-Civita connection:
@@
 In the following we will denote with $\nabla_{\mu}$ the covariant derivative with respect to the Levi-Civita connection $\Gamma_{\mu}{}^{\varrho}{}_{\nu}$, and with $R^{\varrho}{}_{\sigma\mu\nu}$ the Riemann tensor.
-There is an alternative but equivalent formulation of General Relativity (at least classically and without boundary terms), known as teleparallel gravity, in which one chooses a connection $W_{\mu}{}^{\varrho}{}_{\nu}$, compatible with the metric, such that the dynamical variables are encoded in the torsion and the curvature vanishes Einstein1928 ; Weitzenbock1928 :
+There is an alternative but equivalent formulation of General Relativity (at least classically and without boundary terms), known as teleparallel equivalent of General Relativity, in which one chooses a connection $W_{\mu}{}^{\varrho}{}_{\nu}$, compatible with the metric, such that the dynamical variables are encoded in the torsion and the curvature vanishes Einstein1928 ; Weitzenbock1928 :
 $\ignorespaces\nabla_{\varrho}(W)\,g_{\mu\nu}$ $=\partial_{\varrho}\,g_{\mu\nu}-W_{\varrho}{}^{\sigma}{}_{\mu}\,g_{\sigma\nu}-W_{\varrho}{}^{\sigma}{}_{\nu}\,g_{\mu\sigma}=0,$ (3.3a)
-$T_{\mu}{}^{\varrho}{}_{\nu}(W)$ $=W_{\mu}{}^{\varrho}{}_{\nu}-W_{\nu}{}^{\varrho}{}_{\mu},$ (3.3b)
-$R_{\mu\nu}{}^{\varrho}{}_{\sigma}(W)$ $=\partial_{\mu}\,W_{\nu}{}^{\varrho}{}_{\sigma}-\partial_{\nu}\,W_{\mu}{}^{\varrho}{}_{\sigma}+W_{\mu}{}^{\varrho}{}_{\tau}\,W_{\nu}{}^{\tau}{}_{\sigma}-W_{\nu}{}^{\varrho}{}_{\tau}\,W_{\mu}{}^{\tau}{}_{\sigma}=0.$ (3.3c)
-While in the usual formulation of General Relativity the gravitational interaction is due to the curvature of spacetime, in teleparallel gravity the gravitational interaction is mediated by a Lorentz-like force in a flat spacetime, as in Yang-Mills theories deAndrade:1997gka .
+[…]
+[…]
+While in the usual formulation of General Relativity the gravitational interaction is due to the curvature of spacetime, in teleparallel formulation the gravitational interaction is mediated by a Lorentz-like force in a flat spacetime, as in Yang-Mills theories deAndrade:1997gka .
 Indeed, while in the former case the action is linear in the curvature, in the latter one the action is quadratic in the torsion.
-Since the spacetime manifold is supposed to be flat, then it is possible to choose a system of locally inertial coordinates valid everywhere.
-In other words, the frame or inverse vielbein $e^{\mu}{}_{a}$ is a set of $d$ vectors in $d$ dimensions in the tangent space of the manifold such that, when it is evaluated in a point, it forms a basis for the tangent space in that point, for all the points of the manifold.
-In such a case the manifold is said parallelisable, and it is possible to choose a connection such that the vielbein is parallel or covariantly constant
-$\partial_{\mu}\,e_{\nu}{}^{a}-W_{\mu}{}^{\varrho}{}_{\nu}\,e_{\varrho}{}^{a}=0,$ (3.4)
+Denoting with $e_{\mu}{}_{a}$ the components of the vielbein and with $e^{\mu}{}_{a}$ the components of the frame or inverse vielbein, it is possible to choose a connection $W_{\mu}{}^{\varrho}{}_{\nu}$ such that the vielbein is parallel or covariantly constant, that is the following equation holds:
+$\partial_{\mu}\,e_{\nu}{}^{a}-W_{\mu}{}^{\varrho}{}_{\nu}\,e_{\varrho}{}^{a}=0.$ (3.4)
 The unique connection solving the previous condition is the Weitzenböck connection Weitzenbock1928
@@
 [^5] The spin connection corresponding to the Weitzenböck connection is zero (Weitzenböck condition).
-See A for more details.
+See Appendix A for more details.
 Moreover, the condition (3.4) implies the compatibility condition (3.3a), since
@@
 The relation between the Levi-Civita connection and the Weitzenböck is given by:
-$\Gamma_{\mu}{}^{\varrho}{}_{\nu}=W_{\mu}{}^{\varrho}{}_{\nu}+K_{\mu}{}^{\varrho}{}_{\nu}=W_{\mu}{}^{\varrho}{}_{\nu}+\tfrac{1}{2}\,(T_{\mu\nu}{}^{\varrho}+T_{\nu\mu}{}^{\varrho}-T_{\mu}{}^{\varrho}{}_{\nu}),$ (3.11)
-where the tensor $-K_{\mu}{}^{\varrho}{}_{\nu}$ is the contorsion.
+[…]
+where the tensor $K_{\mu}{}^{\varrho}{}_{\nu}$ is the contorsion.
 Using the vanishing of the curvature of the Weitzenböck connection and the relation (3.11), one can write the Riemann tensor in terms of the contorsion in the following way:
-$R^{\mu}{}_{\nu\varrho\sigma}=D_{\varrho}\,K_{\sigma}{}^{\mu}{}_{\nu}-D_{\sigma}\,K_{\rho}{}^{\mu}{}_{\nu}+T_{\varrho}{}^{\tau}{}_{\sigma}\,K_{\tau}{}^{\mu}{}_{\nu}+K_{\varrho}{}^{\mu}{}_{\tau}\,K_{\sigma}{}^{\tau}{}_{\nu}-K_{\sigma}{}^{\mu}{}_{\tau}\,K_{\varrho}{}^{\tau}{}_{\nu}.$ (3.12)
+$R^{\mu}{}_{\nu\varrho\sigma}=D_{\varrho}\,K_{\sigma}{}^{\mu}{}_{\nu}-D_{\sigma}\,K_{\rho}{}^{\mu}{}_{\nu}+T_{\varrho}{}^{\tau}{}_{\sigma}\,K_{\tau}{}^{\mu}{}_{\nu}+K_{\varrho}{}^{\mu}{}_{\tau}\,K_{\sigma}{}^{\tau}{}_{\nu}-K_{\sigma}{}^{\mu}{}_{\tau}\,K_{\varrho}{}^{\tau}{}_{\nu}$.
+(3.12)
 Therefore, using the trace of the Bianchi identity (3.9), the Ricci tensor is
@@
 $+\tfrac{1}{4}\,T_{\varrho\mu\sigma}\,T^{\varrho}{}_{\nu}{}^{\sigma}-\tfrac{1}{2}\,T_{\mu\varrho\sigma}\,T_{\nu}{}^{\varrho\sigma}-\tfrac{1}{2}\,T_{\mu\varrho\sigma}\,T_{\nu}{}^{\sigma\varrho}$, (3.13)
-whose the Ricci scalar is
-$R=R^{\mu}{}_{\mu}=-2\,D_{\mu}\,T^{\mu\nu}{}_{\nu}-\tfrac{1}{4}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}-\tfrac{1}{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}-T_{\mu\nu}{}^{\nu}\,T^{\mu\varrho}{}_{\varrho}.$ (3.14)
+and the Ricci scalar is
+$R=R^{\mu}{}_{\mu}=-2\,D_{\mu}\,T^{\mu\nu}{}_{\nu}-\tfrac{1}{4}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}-\tfrac{1}{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}-T_{\mu\nu}{}^{\nu}\,T^{\mu\varrho}{}_{\varrho}$.
+(3.14)
 Finally, using the Weitzenböck covariant divergence (3.10), one gets the following identity
-$e\,R=-\partial_{\mu}\,(2\,e\,T^{\mu\nu}{}_{\nu})-e\,(\tfrac{1}{4}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\tfrac{1}{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}-T_{\mu\nu}{}^{\nu}\,T^{\mu\varrho}{}_{\varrho}),$ (3.15)
+[…]
 which shows that, up to a boundary term, the Einstein-Hilbert action admits an equivalent formulation in terms of a linear combination of the quadratic scalar contractions of the Weitzenböck torsion Moller1961a ; Moller1961b ; Moller1978 ; Hayashi:1967se ; Hayashi:1979qx ; deAndrade:1997gka :
@@
 where we set $\kappa=8\,\pi\,G=1$.
+[^6] This proves the dynamical equivalence of the usual formulation of General Relativity and of the teleparallel equivalent, since it is not necessary to assume the existence of a globally defined frame Cederwall:2021xqi .
+Although not manifest in teleparallel formulation, the precise combination in (3.16) is locally Lorentz invariant (see next Sections and Cederwall:2021xqi ), and the inverse frame serves as a gauge field.
+Thus, the global extension of the frame is ensured by suitably choosing the gauge redundancy in the overlaps of the coordinates patches.
 4 Review of Møller-Hayashi-Shirafuji theory
 The relation (3.16) shows that the Einstein-Hilbert action consists in a linear combination of three quadratic scalar contractions of the Weitzenböck torsion with precise relative coefficients.
-In Moller1961a ; Moller1961b ; Moller1978 ; Hayashi:1967se ; Hayashi:1979qx the theory defined by the same action with arbitrary coefficients is considered:
-$S_{\textsc{mhs}}=\frac{1}{2}\int\mathrm{d}^{d}x\,e\,(\alpha_{1}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\alpha_{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}+\alpha_{3}\,T^{\mu\nu}{}_{\nu}\,T_{\mu\varrho}{}^{\varrho}).$ (4.1)
+In Moller1961a ; Moller1961b ; Moller1978 ; Hayashi:1967se ; Hayashi:1979qx the theory defined by the same action with arbitrary coefficients is considered,
+[…]
 We will call the theory described by this action Møller-Hayashi-Shirafuji theory (mhs).
@@
 is captured by formally requiring the local Lorentz invariance, which indeed is the necessary requirement in the vielbein formulation of General Relativity, beside diffeomorphism invariance, in order to recover the right number of degrees of freedom Cho:1975dh .
+[^8] Since in mhs theory local Lorentz-invariance is spoiled, unless (4.2) are met, one has to assume the parallelisability of spacetime a priori, that is the existence of a frame which serves as basis for the tangent bundle, which trivialises Hayashi:1979qx .
 The mhs action is the most general quadratic action in the Weitzenböck torsion, since there are only three possible independent quadratic scalar contractions of the torsion.
-Indeed, the torsion has three irreducible components:
-$T_{\mu\varrho\nu}\rightarrow(1,1)\otimes(1)=(1,1,1)\otimes(2,1)_{t}\otimes(1),$ (4.3)
-where $(\lambda_{1},\lambda_{2},\dots,\lambda_{p})$ denotes a Young tableau with $p$ rows and $\lambda_{i}$ boxes in the $i$-th row, $i=1,\dots,p$; and $t$ denotes the traceless part.
+Indeed, the torsion has three irreducible components: [^9] See note 2 for the notation of Young tableaux.
+$T_{\mu\varrho\nu}\rightarrow(1,1)\otimes(1)=(1,1,1)\otimes(2,1)_{t}\otimes(1).$ (4.3)
 $(1,1,1)$ is the totally antisymmetric part $T_{[\mu\nu\varrho]}$, and $(1)$ is the unique trace $T_{\mu\nu}{}^{\nu}$.
@@
 $-2\,T_{\nu\mu\alpha}\,T^{\alpha\beta}{}_{\beta}-2\,D_{\alpha}\,T_{\mu}{}^{\alpha}{}_{\nu}-2\,D_{\alpha}\,T_{\nu\mu}{}^{\alpha}$, (4.6b)
-$\frac{e_{\mu a}}{e}\,\frac{\delta}{\delta e^{\nu}{}_{a}}\,\int\mathrm{d}^{d}x\,e\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}=$ $\,g_{\mu\nu}\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}+2\,g_{\mu\nu}\,D_{\alpha}\,T^{\alpha\beta}{}_{\beta}-2\,D_{\nu}\,T_{\mu\alpha}{}^{\alpha}.$ (4.6c)
+$\frac{e_{\mu a}}{e}\,\frac{\delta}{\delta e^{\nu}{}_{a}}\,\int\mathrm{d}^{d}x\,e\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}$= $\,g_{\mu\nu}\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}+2\,g_{\mu\nu}\,D_{\alpha}\,T^{\alpha\beta}{}_{\beta}-2\,D_{\nu}\,T_{\mu\alpha}{}^{\alpha}$.
+(4.6c)
 Therefore, the equations of motion $E_{\mu\nu}:=\frac{e_{\mu a}}{e}\frac{\delta S_{\textsc{mhs}}}{\delta e^{\nu}{}_{a}}=0$, decomposed in symmetric and antisymmetric part $E_{\mu\nu}=E_{\mu\nu}^{(\text{sym})}+E_{\mu\nu}^{\text{(anti)}}$, with $E_{\mu\nu}^{\text{(sym)}}=E_{\nu\mu}^{\text{(sym)}}$, and $E_{\mu\nu}^{\text{(anti)}}=-E_{\nu\mu}^{\text{(anti)}}$, respectively read
@@
 $-(\alpha_{1}+\tfrac{\alpha_{2}}{2})(T_{(\mu\nu)}{}^{\alpha}\,T_{\alpha\beta}{}^{\beta}+D_{\alpha}\,T_{(\mu\nu)}{}^{\alpha})-\tfrac{\alpha_{3}}{2}\,D_{(\mu}\,T_{\nu)\alpha}{}^{\alpha}\,$+
-$-\tfrac{1}{2}\,g_{\mu\nu}\,(\alpha_{1}\,T^{\alpha\beta\gamma}\,T_{\alpha\beta\gamma}+\alpha_{2}\,T^{\alpha\beta\gamma}\,T_{\alpha\gamma\beta}-\alpha_{3}\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}-2\,\alpha_{3}\,D_{\alpha}\,T^{\alpha\beta}{}_{\beta}),$ (4.7a)
-$E_{\mu\nu}^{\text{(anti)}}=$ $\,(\tfrac{\alpha_{2}}{2}-\alpha_{1})\,(T_{[\mu\nu]}{}^{\alpha}\,T_{\alpha\beta}{}^{\beta}+D_{\alpha}\,T_{[\mu\nu]}{}^{\alpha})-(\alpha_{2}+\tfrac{\alpha_{3}}{2})\,(T_{\mu\alpha\nu}\,T^{\alpha\beta}{}_{\beta}+D_{\alpha}\,T_{\mu}{}^{\alpha}{}_{\nu}).$ (4.7b)
+[…]
+[…]
 where in the antisymmetric part we use the trace of the Bianchi identity (3.9).
@@
 $+\tfrac{1}{2}\,T_{\alpha\mu\beta}\,T^{\alpha}{}_{\nu}{}^{\beta}+T_{(\mu\nu)}{}^{\alpha}\,T_{\alpha\beta}{}^{\beta}+D_{\alpha}\,T_{(\mu\nu)}{}^{\alpha}$, (4.8)
-$D_{\alpha}\,T^{\alpha\beta}{}_{\beta}=$ $\,-\tfrac{1}{2}\,R-\tfrac{1}{8}\,T^{\alpha\beta\gamma}\,T_{\alpha\beta\gamma}-\tfrac{1}{4}\,T^{\alpha\beta\gamma}\,T_{\alpha\gamma\beta}-\tfrac{1}{2}\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}.$ (4.9)
+$D_{\alpha}\,T^{\alpha\beta}{}_{\beta}$= $\,-\tfrac{1}{2}\,R-\tfrac{1}{8}\,T^{\alpha\beta\gamma}\,T_{\alpha\beta\gamma}-\tfrac{1}{4}\,T^{\alpha\beta\gamma}\,T_{\alpha\gamma\beta}-\tfrac{1}{2}\,T^{\alpha\beta}{}_{\beta}\,T_{\alpha\gamma}{}^{\gamma}$.
+(4.9)
 Moreover, it will be useful for the following to decompose the torsion in the components in transforming in the Young tableaux $(2,1)$ and $(1,1,1)$, the first one including the trace:
 $T_{\mu\nu\varrho}=f_{\mu\varrho\nu}+t_{\mu\nu\varrho},$ (4.10)
-where [^7] The inverse relations read: $f_{\mu\nu\varrho}=\tfrac{1}{3}\,(T_{\mu\nu\varrho}+T_{\varrho\mu\nu}-2\,T_{\nu\varrho\mu})$ and $t_{\mu\nu\varrho}=\tfrac{1}{3}\,(T_{\mu\nu\varrho}+T_{\varrho\mu\nu}+T_{\nu\varrho\mu})$.
+where [^10] The inverse relations read $f_{\mu\nu\varrho}=\tfrac{1}{3}\,(T_{\mu\nu\varrho}+T_{\varrho\mu\nu}-2\,T_{\nu\varrho\mu})$ and $t_{\mu\nu\varrho}=\tfrac{1}{3}\,(T_{\mu\nu\varrho}+T_{\varrho\mu\nu}+T_{\nu\varrho\mu})$.
 $f_{\mu\nu\varrho}=-f_{\nu\mu\varrho},\quad f_{[\mu\nu\varrho]}=0,\quad t_{\mu\nu\varrho}=\tfrac{1}{6}\,t_{[\mu\nu\varrho]}.$ (4.11)
@@
 $\mathring{E}_{\mu\nu}^{\text{(sym)}}=$ $\,\alpha_{3}\,(\mathring{R}_{\mu\nu}-\tfrac{1}{2}\,\mathring{R}\,\eta_{\mu\nu})+\tfrac{1}{2}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\partial_{\alpha}\,\mathring{f}^{\alpha}{}_{(\mu\nu)}$, (5.5)
-$\mathring{E}_{\mu\nu}^{\text{(anti)}}=$ $\,(-2\,\alpha_{1}+2\,\alpha_{2}+\tfrac{\alpha_{3}}{2})\,\partial_{\alpha}\,\mathring{t}^{\alpha}{}_{\mu\nu}+\tfrac{1}{2}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\partial_{\alpha}\,\mathring{f}^{\alpha}{}_{[\mu\nu]}.$ (5.6)
+$\mathring{E}_{\mu\nu}^{\text{(anti)}}=$ $\,(-2\,\alpha_{1}+2\,\alpha_{2}+\tfrac{\alpha_{3}}{2})\,\partial_{\alpha}\,\mathring{t}^{\alpha}{}_{\mu\nu}+\tfrac{1}{2}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\partial_{\alpha}\,\mathring{f}^{\alpha}{}_{[\mu\nu]}$.
+(5.6)
 where
@@
 Since the mhs theory is formulated in a quadratic way, unlike General Relativity in conventional formulation, the linearised equations of motion can be obtained directly by varying the action (4.1) with the torsion $T_{\mu\nu\varrho}$ replaced by the linearised one $\mathring{T}_{\mu\nu\varrho}$:
-$\frac{1}{2}\int\mathrm{d}^{d}\,x\,(\alpha_{1}\,\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\nu\varrho}+\alpha_{2}\,\mathring{T}^{\mu\nu\varrho}\,\mathring{T}_{\mu\varrho\nu}+\alpha_{3}\,\mathring{T}^{\mu\nu}{}_{\nu}\,\mathring{T}_{\mu\varrho}{}^{\varrho}).$ (5.13)
+[…]
 This action can be obtained working directly at the linear level in the following way.
@@
 $s\,h_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}+\partial_{\nu}\,\mathring{\xi}_{\mu},\quad s\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},\quad\quad s\,\mathring{\xi}^{\mu}=0.$ (5.15)
-So, one has to study the local $s$-cohomology of ghost number zero on the jet space of $\overline{h}_{\mu\nu}$ and its derivatives.
+So, one has to study the local $s$-cohomology of ghost number zero on the space of $\overline{h}_{\mu\nu}$ and its derivatives, treated as independent variables (the so-called jet space).
 This is done in Appendix B.
@@
 On the other hand, the transformation of the symmetric part is the same as before $s_{\textsc{fp}}\,h_{\mu\nu}=s\,h_{\mu\nu}$.
-Therefore, the local $s_{\textsc{fp}}$-cohomology on the jet space of $\{h_{\mu\nu},\mathring{\xi}_{\mu},b_{\mu\nu},\tilde{\Omega}_{\mu\nu}\}$ and their derivatives is equivalent to the local $s$-cohomology on the jet space of the starting field space $\{h_{\mu\nu},\mathring{\xi}_{\mu}\}$.
+Therefore, the local $s_{\textsc{fp}}$-cohomology on the space of $\{h_{\mu\nu},\mathring{\xi}_{\mu},b_{\mu\nu},\tilde{\Omega}_{\mu\nu}\}$ and their derivatives is equivalent to the local $s$-cohomology on the space of the starting field space $\{h_{\mu\nu},\mathring{\xi}_{\mu}\}$ and derivatives.
 The linearised torsion $\mathring{T}_{\mu\nu\varrho}$ is not invariant under the modified transformations, but it transforms with respect to $\mathring{\Omega}_{\mu\nu}$ as a gauge two-form with a spectator index
@@
 Consistently, the action (5.13) with coefficients (4.2) does not depend on the antisymmetric part $b_{\mu\nu}$ by integrating by parts.
-6 Embedding covariant fractons in linearised MHS theory
-In this section we show that all the solutions of the equations of motion of covariant fracton gauge theory (2.5) or (2.9) are solutions of the equations of motion of linearised mhs theory (5.5)–(5.6).
-The space of the solutions of covariant fracton gauge theory modulo the gauge invariance (fracton moduli space) is the submanifold $\mathscr{M}_{\text{fr}}$ in the space of fields identified by the equations of motion (2.5) or (2.9),
-$\mathscr{M}_{\text{fr}}=\{f^{\text{(fr)}}_{\mu\nu\varrho}(x)\,:\,E^{\text{(fr)}}_{\mu\nu}(f^{\text{(fr)}})=0\}$ (6.1)
-where the elements in $\mathscr{M}_{\text{fr}}$ are identified by the invariant fracton field strength because of the gauge invariance of the equations of motion: two solutions $h^{\text{(fr)}}_{\mu\nu}(x)$ and $\tilde{h}^{\text{(fr)}}_{\mu\nu}(x)$ which differ by a fracton gauge transformation are identified in $\mathscr{M}_{\text{fr}}$.
-Similarly, the moduli space of the linearised mhs theory is the submanifold $\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}$ of the field space identified by the equations of motion (5.5)–(5.6):
-$\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}=\{\mathring{T}_{\mu\nu\varrho}(x)\,:\,\mathring{E}^{\text{(sym)}}_{\mu\nu}(\mathring{T})=0,\;\mathring{E}^{\text{(anti)}}_{\mu\nu}(\mathring{T})=0\}.$ (6.2)
-Consider an arbitrary fixed element $f_{\mu\nu\varrho}^{\text{(fr)}}$ in $\mathscr{M}_{\text{lin}}$. If we consider a linearised torsion $\tilde{T}_{\mu\nu\varrho}$, such that
-$\tilde{T}_{\mu\nu\varrho}=\tfrac{1}{2}\,f_{\mu\varrho\nu}^{\text{(fr)}},$ (6.3)
-which means, in particular, that we are considering a linearised torsion with vanishing totally antisymmetric part, then $\tilde{T}_{\mu\nu\varrho}$ is an element in $\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}$.
-To see why, consider that
-$\ignorespaces\mathring{E}_{\mu\nu}^{\text{(sym)}}(\tilde{T})$ $=\alpha_{3}\,(R^{(\text{fr})}_{\mu\nu}-\tfrac{1}{2}\,R^{(\text{fr})}\,\eta_{\mu\nu})+\tfrac{1}{4}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\partial_{\alpha}\,f^{(\text{fr})\,\alpha}{}_{(\mu\nu)},$ (6.4a)
-$\mathring{E}_{\mu\nu}^{\text{(anti)}}(\tilde{T})$ $=\tfrac{1}{4}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\partial_{\alpha}\,f^{(\text{fr})\,\alpha}{}_{[\mu\nu]},$ (6.4b)
-where $R^{(\text{fr})}_{\mu\nu}$ and $R^{(\text{fr})}$ are given by (2.7) and (2.8).
-Notice that the constants $\alpha_{1}$ and $\alpha_{2}$ appear only in the combination $2\,\alpha_{1}+\alpha_{2}$, which means that, had we varied the action with, say, $\alpha_{2}=0$, we would have obtained the same result.
+6 Particle spectrum in covariant fracton gauge theory
+The space of the solutions of covariant fracton gauge theory $\mathscr{M}_{\text{fr}}$ (moduli space) is the submanifold in the space of fields $\{h^{\text{(fr)}}_{\mu\nu}\}$, identified by the equations (2.5) or (2.10), modulo the gauge invariance, [^12] The author thanks C.
+Imbimbo, who suggested how to expand this and the next section which, as stated by the brst rule in (2.1), [^13] In this section $\lambda$ is a simple gauge parameter.
+We maintain the same notation of the corresponding ghost because this should not cause confusion.
+amounts to linearised longitudinal diffeomorphisms,
+$\delta_{\text{fracton}}\,h_{\mu\nu}^{\text{(fr)}}=\partial_{\mu}\,\partial_{\nu}\,\lambda,\;\;\forall\;\alpha,\beta.$ (6.1)
+This means that two solutions $h^{\text{(fr)}}_{\mu\nu}(x)$ and $h^{\prime\text{(fr)}}_{\mu\nu}(x)$ of the equations of motion, which differ by a fracton gauge transformation are identified in the moduli space.
+When $\frac{\beta}{\alpha}=-2$, as in (2.11), the theory reduced to linearised Einstein gravity, and the gauge invariance is extended to all linearised diffeomorphisms:
+$\delta_{\text{diff}}\,h_{\mu\nu}^{\text{(fr)}}=\partial_{\mu}\,\xi_{\nu}+\partial_{\nu}\,\xi_{\mu},\;\;\text{if}\;\;\tfrac{\beta}{\alpha}=-2.$ (6.2)
+There is also a case in which the fracton theory is traceless, that is, it does not depend on $h^{\text{(fr)}\varrho}{}_{\varrho}$, when $h_{\mu\nu}^{\text{(fr)}}$ is decomposed in its traceless part and in its trace.
+One can check that this happens if and only if
+$\tfrac{\beta}{\alpha}=-\tfrac{2}{d-1},$ (6.3)
+where $\alpha$ and $\beta$ are the parameters introduced in (2).
+In that case, there is an additional gauge symmetry, whose effect is to shift the trace, corresponding to a linearised Weyl scaling of $h^{\text{(fr)}}_{\mu\nu}$:
+$\delta_{\text{Weyl}}\,h^{\text{(fr)}}_{\mu\nu}=2\,\sigma\,\eta_{\mu\nu},\;\;\text{if}\;\;\tfrac{\beta}{\alpha}=-\tfrac{2}{d-1}.$ (6.4)
+Let us find the general solution of covariant fracton gauge theory.
+For definiteness, we consider the four dimensional case from now on.
+The problem was addressed in Afxonidis:2023pdq ; Afxonidis:2024tph , where the particle content of the theory was studied: except for some special cases, the theory describes five propagating degrees of freedom, with helicities $0,\pm 1,\pm 2$ respectively.
+Working in Fourier space, $p^{\mu}$ denotes the momentum and $\tilde{h}^{\text{(fr)}}_{\mu\nu}$ the Fourier transform of the $h^{\text{(fr)}}_{\mu\nu}$.
+The equations (2.10) in Fourier space are:
+$K_{\mu\nu}{}^{\varrho\sigma}(p^{2};\alpha,\beta)\,\tilde{h}^{\text{(fr)}}_{\varrho\sigma}$ $=2\,\alpha\,p^{2}\,\tilde{h}^{\text{(fr)}}_{\mu\nu}-\tfrac{1}{2}\,(2\,\alpha-\beta)\,p_{\alpha}\,p_{(\mu}\,\tilde{h}^{\text{(fr)}}_{\nu)}{}^{\alpha}\,$+
+$-\beta\,p_{\mu}\,p_{\nu}\,\tilde{h}^{\text{(fr)}}+\beta\,\eta_{\mu\nu}\,(p^{2}\,\tilde{h}^{\text{(fr)}}-p_{\alpha}\,p_{\beta}\,\tilde{h}^{\text{(fr)}\,\alpha\beta})=0,$ (6.5)
+where $K(p^{2};\alpha,\beta)$, viewed as a $10\times 10$ matrix, identifies a system of homogeneous linear equations in the $10$ variables $\{\tilde{h}^{\text{(fr)}}_{\mu\nu}\}$.
+Trivial or pure-gauge solutions are those which solve the system without any condition on the momentum, and they correspond to the gauge invariance of the theory.
+The number of independent trivial solutions is given by the dimension of the kernel of off-shell $K(p^{2};\alpha,\beta)$ (null eigenspace, whose dimension is the number of variables decreased by the matrix rank).
+One can show that
+$\text{trivial solutions}=\dim\,\text{ker}\,K(p^{2}\neq 0;\alpha,\beta)=\begin{cases}6&\text{if}\;\alpha=0,\\ 4&\text{if}\;\tfrac{\beta}{\alpha}=-2,\\ 2&\text{if}\;\tfrac{\beta}{\alpha}=-\tfrac{2}{3},\\ 1&\text{otherwise}.\end{cases}$ (6.6)
+$p_{\mu}\,p_{\nu}\,\lambda$ is always a trivial solution; in the linearised Einstein case $\frac{\beta}{\alpha}=-2$, this solution is encompassed in $p_{(\mu}\,\xi_{\nu)}$, which are four independent trivial solutions; in the traceless case $\frac{\beta}{\alpha}=-\frac{2}{3}$, besides $p_{\mu}\,p_{\nu}\,\lambda$, $2\,\eta_{\mu\nu}\,\sigma$ is trivial too.
+There is one peculiar case more, $\alpha=0$, in which there is no kinetic term for the fracton gauge field, so that the theory is expected to have no propagating degrees of freedom in this case Blasi:2022mbl .
+In the massless case, the momentum can be canonically parametrised as $p^{\mu}=(1,0,0,1)$, such that $p^{2}=0$.
+The dimension of the kernel in the massless case gives the number of independent solutions without taking into account the gauge redundancy.
+The result is
+$\dim\,\text{ker}\,K(p^{2}=0,\alpha,\beta)=\begin{cases}8\quad\text{if}\;\frac{\beta}{\alpha}=2,\\ 6\quad\text{otherwise}.\end{cases}$ (6.7)
+The case $\frac{\beta}{\alpha}=2$ is special, because the contraction $p^{\alpha}\,\tilde{h}^{(\text{fr})}_{\alpha\mu}$ disappears from the equations (6.5).
+Subtracting the number of trivial solutions to the dimension of the kernel in the various cases, one gets the number of degrees of freedom of the theory:
+$_{\text{dof}}=\begin{cases}6-6=0,&\text{if}\;\alpha=0,\\ 6-4=2,&\text{if}\;\tfrac{\beta}{\alpha}=-2\;\text{(linearised gravity)},\\ 6-2=4,&\text{if}\;\tfrac{\beta}{\alpha}=-\tfrac{2}{3}\;\text{(traceless limit)},\\ 8-1=7,&\text{if}\;\tfrac{\beta}{\alpha}=2,\\ 6-1=5,&\text{otherwise,}\end{cases}$ (6.8)
+where we see that, as expected, the theory has no propagating degrees of freedom when $\alpha=0$.
+In the case of linearised Einstein gravity, the two degrees of freedom must describe the graviton, with helicities $\pm 2$.
+In the third case, we expect that the tracelessness condition eliminates a scalar particle, so that in the general case, which has five degrees of freedom, at least a scalar particle is included in the spectrum.
+Since the general case encompasses also the second, the graviton is also included.
+Only two degrees of freedom remain to be identified: they turn out to describe a particle with helicities $\pm 1$, as confirmed by the explicit solutions of the equations of motion.
+The two degrees of freedom more in the special case $\tfrac{\beta}{\alpha}=2$ have again helicities $\pm 1$, as we will see in the end of this section.
+Solving the equations of motion means finding a basis for the kernel in the massless case, eliminating the trivial solutions.
+In the general case, apart from the trivial solutions $p_{\mu}\,p_{\nu}\,\lambda$, the other five independent elements in the basis can be chosen to be equal to
+$\text{diag}\,(-1,-\tfrac{2\,\alpha+\beta}{2\,\beta},-\tfrac{2\,\alpha+\beta}{2\,\beta},1),\;\;\lambda_{\mu\nu}^{\pm 1},\;\;\lambda_{\mu\nu}^{\pm 2},\;\;\text{if}\;\beta\neq 0,\alpha\neq 0,\tfrac{\beta}{\alpha}=\pm 2,-\tfrac{2}{3},$ (6.9)
+where $\lambda_{\mu\nu}^{s}$ are polarisation matrices with helicity $s=\pm 1,\pm 2$ (the explicit expressions are in (D.4b)–(D.4c)), and the diagonal matrix has vanishing helicity.
+Therefore, in the general case, covariant fracton gauge theory describes particles with helicities $0,\pm 1,\pm 2$ Afxonidis:2023pdq .
+If $\beta=0$, the scalar solution in (6.9) is replaced by
+$\text{diag}\,(0,1,1,0),$ (6.10)
+which again has vanishing helicity.
+If $\frac{\beta}{\alpha}=-2$, the trivial solutions $p_{(\mu}\,\xi_{\nu)}$ are completed to a basis for the kernel by $\lambda_{\mu\nu}^{\pm 2}$, as expected in linearised gravity.
+If $\frac{\beta}{\alpha}=-\frac{2}{3}$, the scalar solution in (6.9) becomes $\eta_{\mu\nu}$, since $\frac{2\,\alpha+\beta}{2\,\beta}\rightarrow 1$, so it is trivial, and the theory describes only the modes $\pm 1,\pm 2$.
+Finally, in the special case $\frac{\beta}{\alpha}=2$, the basis has eight independent elements: one of them is the trivial one $p_{\mu}\,p_{\nu}\,\lambda$; the remaining seven ones can be chosen as
+$\text{diag}\,(-1,-1,-1,1),\;\;\begin{pmatrix}0&-1&\mp i&0\\ -1&0&0&0\\ \mp i&0&0&0\\ 0&0&0&0\end{pmatrix},\;\;\begin{pmatrix}0&0&0&0\\ 0&0&0&-1\\ 0&0&0&\mp i\\ 0&-1&\mp i&0\end{pmatrix},\;\;\lambda_{\mu\nu}^{\pm 2},$ (6.11)
+where the diagonal solution has vanishing helicity, $\lambda_{\mu\nu}^{\pm 2}$ has helicities $\pm 2$, and the remaining two pairs of solutions have both helicities $\pm 1$.
+So, compared to the general case, there is one $\pm 1$ helicity particle more.
+As a last observation, consider that, the polarisation matrices $\lambda_{\mu\nu}^{s}$ (D.4a)–(D.4c), in the general solution (6.9), are transversal in the massless case, meaning that
+$p^{\mu}\,\lambda_{\mu\nu}^{s}=0,\;\;s=\pm 1,\pm 2,\;\;\text{if}\;p^{\mu}=(1,0,0,1).$ (6.12)
+Instead, the scalar mode is longitudinal
+$p^{\mu}\,\text{diag}\,(-1,-\tfrac{2\,\alpha+\beta}{2\,\beta},-\tfrac{2\,\alpha+\beta}{2\,\beta},1)_{\mu\nu}=p_{\nu}.$ (6.13)
+Nonetheless, this implies that the arbitrary solution $\tilde{h}_{\mu\nu}^{(\text{fr})}$ on-shell satisfies the condition
+$p^{\alpha}\,p_{[\mu}\,\tilde{h}^{\text{(fr)}}_{\nu]\alpha}=0\;\;\text{if}\;p^{\mu}=(1,0,0,1),$ (6.14)
+which will be useful in the following.
+7 Particle spectrum of MHS theory and fracton embedding
+Let us now find the particle spectrum for the mhs theory (see also Hayashi:1979qx ), following the same method as in the previous section.
+It will be convenient to redefine the constants $\alpha_{1}$ and $\alpha_{2}$ as
+$\tilde{\alpha}_{1}=2\,\alpha_{1}+\alpha_{2},\quad\tilde{\alpha}_{2}=2\,\alpha_{1}-\alpha_{2}.$ (7.1)
+The moduli space $\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}$ of the linearised mhs theory is the submanifold in the field space $\{h_{\mu\nu}$, $b_{\mu\nu}\}$ identified by the equations of motion (5.5)–(5.6), modulo the gauge invariance, which amounts to linearised diffeomorphisms acting both on $h_{\mu\nu}$ and $b_{\mu\nu}$, [^14] In this section $\xi_{\mu}$, and $\Omega_{\mu\nu}$ are simple gauge parameters.
+We maintain the same notation of the corresponding ghosts (omitting the dot denoting the linearisation) because this should not be cause confusion.
+$\delta_{\text{diff}}\,h_{\mu\nu}=\partial_{\mu}\,\xi_{\nu}+\partial_{\nu}\,\xi_{\mu},\;\delta_{\text{diff}}\,b_{\mu\nu}=\partial_{\mu}\,\xi_{\nu}-\partial_{\nu}\,\xi_{\mu},\;\;\forall\;\tilde{\alpha}_{1},\tilde{\alpha}_{2},\alpha_{3},$ (7.2)
+as in (5.15).
+As in covariant fracton gauge theory, there are some particular cases in which the gauge symmetry is enlarged.
+Linearised Einstein gravity is reproduced when the parameters are as in (4.2), or equivalently $\tilde{\alpha}_{2}=0,\frac{\alpha_{3}}{\tilde{\alpha}_{1}}=-1$, according to (7.1), up to an overall constant.
+In this case there is no dependence on $b_{\mu\nu}$, so it can be shifted away, and this is the effect of local Lorentz invariance, as previously discussed:
+$\delta_{\text{Lorentz}}\,h_{\mu\nu}=0,\;\delta_{\text{Lorentz}}\,b_{\mu\nu}=\Omega_{\mu\nu},\;\;\text{if}\;\;\tilde{\alpha}_{2}=0,\;\tfrac{\alpha_{3}}{\tilde{\alpha}_{1}}=-1.$ (7.3)
+The traceless limit in mhs theory is reached if and only if
+$\tfrac{\alpha_{3}}{\tilde{\alpha}_{1}}=-\tfrac{1}{d-1}.$ (7.4)
+in which case there is no dependence on the trace $h^{\varrho}{}_{\varrho}$ and the gauge invariance includes a Weyl scaling:
+$\delta_{\text{Weyl}}\,h_{\mu\nu}=2\,\sigma\,\eta_{\mu\nu},\;\;\delta_{\text{Weyl}}\,b_{\mu\nu}=0,\;\;\text{if}\;\;\tfrac{\alpha_{3}}{\tilde{\alpha}_{1}}=-\tfrac{1}{d-1}.$ (7.5)
+A last limit case is when the theory does not depend on $h_{\mu\nu}$ at all.
+This happens when $\tilde{\alpha}_{1}=\alpha_{3}=0$, in which case the symmetric equation (5.5) trivialises, and mhs theory reduces to the theory for a free two-form gauge field (Ramond-Kalb field rk Kalb:1974yc ):
+$\delta_{\textsc{rk}}\,h_{\mu\nu}=\Delta_{\mu\nu},\;\delta_{\textsc{rk}}\,b_{\mu\nu}=\partial_{\mu}\,\xi_{\nu}-\partial_{\nu}\,\xi_{\mu},\;\;\text{if}\;\;\tilde{\alpha}_{1}=\alpha_{3}=0,$ (7.6)
+where $\Delta_{\mu\nu}$ is any symmetric rank-two tensor, shifting $h_{\mu\nu}$.
+In order to find the general solution of (5.5)–(5.6), we write them in Fourier space. Again, we consider from now on the four-dimensional case. The equations read
+$\tfrac{1}{2}\,\tilde{\alpha}_{1}\,p^{2}\,\tilde{h}_{\mu\nu}-\tfrac{1}{4}\,(\tilde{\alpha}_{1}-\alpha_{3})\,p_{\alpha}\,p_{(\mu}\,\tilde{h}_{\nu)}{}^{\alpha}-\tfrac{1}{2}\,\alpha_{3}\,p_{\mu}\,p_{\nu}\,\tilde{h}\,$+
+$\quad+\tfrac{1}{2}\,\alpha_{3}\,\eta_{\mu\nu}\,(p^{2}\,\tilde{h}-p_{\alpha}\,p_{\beta}\,\tilde{h}^{\alpha\beta})+\tfrac{1}{4}\,(\tilde{\alpha}_{1}+\alpha_{3})\,p_{\alpha}\,p_{(\mu}\,\tilde{b}_{\nu)}{}^{\alpha}=0,$ (7.7a)
+$\tfrac{1}{2}\,\tilde{\alpha}_{2}\,p^{2}\,\tilde{b}_{\mu\nu}-\tfrac{1}{4}\,(\tilde{\alpha}_{1}-2\,\tilde{\alpha}_{2}+\tilde{\alpha}_{3})\,p^{\alpha}\,p_{[\mu}\,\tilde{b}_{\nu]\alpha}-\tfrac{1}{4}\,(\tilde{\alpha}_{1}+\alpha_{3})\,p^{\alpha}\,p_{[\mu}\,\tilde{h}_{\nu]\alpha}=0.$ (7.7b)
+This is a system of homogeneous linear equations in the $10+6$ variables $(\tilde{h}_{\mu\nu},\tilde{b}_{\mu\nu})$.
+Looking at the equations alone it is possible to deduce the embedding of the fracton moduli space in the mhs one.
+Indeed, one sees that, if [^15] In direct space, this means that the linearised torsion $\mathring{T}_{\mu\nu\varrho}$, corresponding to the choice (7.8), $\mathring{T}_{\mu\nu\varrho}=\tfrac{1}{2}\,f_{\mu\varrho\nu}^{\text{(fr)}}$.
+Moreover, there is no dependence in the projected equations on $\tilde{\alpha}_{2}$, and that the old constants $\alpha_{1}$ and $\alpha_{2}$ appear only in the combination $2\,\alpha_{1}+\alpha_{2}=\tilde{\alpha}_{1}$, which means that, had we varied the action with, say, $\alpha_{2}=0$, we would have obtained the same result.
 This is not surprising, since there are only two independent quadratic scalar contractions if the totally antisymmetric part of the torsion is vanishing.
-The projected symmetric equation (6.4a) corresponds to the equations of motion in covariant fracton gauge theory (2.9)
-$\mathring{E}_{\mu\nu}^{\text{(sym)}}(\tilde{T})=E_{\mu\nu}^{\text{(fr)}}(f^{\text{(fr)}})=0,$ (6.5)
-upon renaming the constant coefficients
-$\alpha=\tfrac{1}{4}\,(2\,\alpha_{1}+\alpha_{2}),\quad\beta=\tfrac{\alpha_{3}}{2}.$ (6.6)
-Nevertheless, in linearised mhs theory there is also the variation with respect to the antisymmetric part of the linearised vielbein, which is responsible for the antisymmetric part of the equations of motion (6.4b).
-But, excluding the case (4.2), which brings to linearised General Relativity and which is such that $02\,\alpha_{1}+\alpha_{2}+\alpha_{3}=$ 0, (6.4b) is implied by (6.4a).
-To show why, consider the first divergence the fracton equations of motion (2.12).
-If we define the following one-form Blasi:2022mbl
-$j=j_{\mu}\,\mathrm{d}x^{\mu}=\partial^{\nu}\,h^{\text{(fr)}}_{\nu\mu}\,\mathrm{d}x^{\mu},$ (6.7)
-the equation (2.12) can be written in this form
-$\partial^{\mu}\,E_{\mu\nu}^{\text{(fr)}}\,\mathrm{d}x^{\nu}=-\mathrm{d}^{\dagger}\,\mathrm{d}\,j=0,$ (6.8)
-where $\mathrm{d}$ and $\mathrm{d}^{\dagger}$ are the De Rham differential and codifferential, as it follows by explicit evaluation
-$\partial^{\mu}\,\partial^{\varrho}\,f^{\text{(fr)}}_{\varrho(\mu\nu)}\,\mathrm{d}x^{\nu}$ $=2\,\partial^{\mu}\,\partial^{\varrho}\,\partial_{\varrho}\,h^{\text{(fr)}}_{\mu\nu}\,\mathrm{d}x^{\nu}-\partial^{\mu}\,\partial^{\varrho}\,\partial_{\mu}\,h^{\text{(fr)}}_{\varrho\nu}\,\mathrm{d}x^{\nu}-\partial^{\mu}\,\partial^{\varrho}\,\partial_{\nu}\,h^{\text{(fr)}}_{\varrho\mu}\,\mathrm{d}x^{\nu}=$
-$=\,\partial^{2}\,\partial^{\mu}\,h^{\text{(fr)}}_{\mu\nu}\,\mathrm{d}x^{\nu}-\mathrm{d}x^{\nu}\,\partial_{\nu}\,\partial^{\mu}\,\partial^{\varrho}\,h^{\text{(fr)}}_{\mu\varrho}=$
-$=\partial^{2}\,j-\mathrm{d}\,\partial_{\mu}\,j^{\mu}=-(\mathrm{d}^{\dagger}\,\mathrm{d}+\mathrm{d}\,\mathrm{d}^{\dagger})\,j+\mathrm{d}\,\mathrm{d}^{\dagger}\,j=-\mathrm{d}^{\dagger}\,\mathrm{d}\,j.$ (6.9)
-Now, the equation (6.8) is equivalent to say that $j$ is $\mathrm{d}$-closed,
-$\partial^{\mu}\,E_{\mu\nu}^{\text{(fr)}}\,\mathrm{d}x^{\nu}=-\mathrm{d}^{\dagger}\,\mathrm{d}\,j=0\Leftrightarrow\mathrm{d}\,j=0,$ (6.10)
-[…]
-$\mathring{E}_{\mu\nu}^{\text{(anti)}}(\tilde{T})\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}=-\tfrac{1}{2}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\mathrm{d}\,j,$ (6.11)
-since
-$\partial^{\varrho}\,f^{\text{(fr)}}_{\varrho\mu\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ $=\partial^{\varrho}\,\partial_{\varrho}\,h^{\text{(fr)}}_{\mu\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}-\partial^{\varrho}\,\partial_{\mu}\,h^{\text{(fr)}}_{\varrho\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}=-\mathrm{d}\,j.$ (6.12)
-In conclusion, we have shown that
-$\mathring{E}_{\mu\nu}^{\text{(anti)}}(\tilde{T})=0\Leftrightarrow\partial^{\mu}\,E_{\mu\nu}^{\text{(fr)}}(f^{(\text{fr})})=0,$ (6.13)
-which means that the projected symmetric equation (6.4a) implies the antisymmetric one (6.4b).
-As a consequence, all the solutions of the covariant fracton gauge theory are solutions of the linearised mhs theory:
-$\mathscr{M}_{\text{fr}}\subset\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}.$ (6.14)
+$\tilde{h}_{\mu\nu}=\tilde{h}_{\mu\nu}^{\text{(fr)}},\quad\tilde{b}_{\mu\nu}=0,\quad\tilde{\alpha}_{1}=4\,\alpha,\quad\alpha_{3}=2\,\beta,$ (7.8)
+then the first equation (7.7a) reduces to the fracton equation (6.5), and the second equation (7.7b) reduces to the condition (6.14).
+This shows that the fracton moduli space, if $\frac{\beta}{\alpha}\neq 2$, is contained in the subsector with vanishing $b_{\mu\nu}$ of the linearised mhs moduli space, upon using the identification in (7.8),
+$\mathscr{M}_{\text{fr}}(\tfrac{\beta}{\alpha}\neq 2)\subseteq{\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}}_{\big|_{b=0}}\subset\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}.$ (7.9)
+Observe that the condition $b_{\mu\nu}=0$ in (7.8) breaks the gauge invariance (7.2), but it leaves a residual gauge freedom when $\xi_{\mu}$ is exact, that is, when it generates longitudinal diffeomorphisms $\xi_{\mu}\propto\partial_{\mu}\,\lambda$.
+This is consistently the fracton gauge invariance (6.1).
+Actually, we can prove an even stronger statement: not only the fracton moduli space is included in the $b_{\mu\nu}=0$ subsector, but it is also isomorphic to the same subsector.
+To see why, let us find the most general solution of the full equations (7.7a)–(7.7b).
+At the very end, we will see that, upon restricting in the $b_{\mu\nu}=0$ subsector, the most general solutions coincides with the fracton one, if $\frac{\beta}{\alpha}\neq 2$.
+A $16\times 16$ matrix $K(p;\tilde{\alpha}_{1},\tilde{\alpha}_{2},\alpha_{3})$ can be introduced, describing the system (7.7a)–(7.7b) in the variables $(\tilde{h}_{\mu\nu},\tilde{b}_{\mu\nu})$, viewed as a vector with $10+6$ entries.
+Except for the special cases in which the gauge invariance is enlarged, the dimension of the kernel in the massless case is 10, and the number of trivial solutions is 4, encoded in $(\tilde{h}_{\mu\nu},\tilde{b}_{\mu\nu})=(p_{(\mu}\,\xi_{\nu)},p_{[\mu}\,\xi_{\nu]})$.
+So, there are six degrees of freedom, which can be chosen to be equal to
+$(\text{diag}\,(-1,-\tfrac{\tilde{\alpha}_{1}+\alpha_{3}}{2\,\alpha_{3}},-\tfrac{\tilde{\alpha}_{1}+\alpha_{3}}{2\,\alpha_{3}},1),O),\;\;(\lambda_{\mu\nu}^{\pm 1},O),\;\;(\lambda_{\mu\nu}^{\pm 2},O),\;\;(O,\lambda_{\mu\nu}^{0}),$ (7.10)
+where $O$ is the $4\times 4$ null matrix, and $\lambda_{\mu\nu}^{s}$ are the polarisation matrices in (D.4a)–(D.4c), with helicities $s=0,\pm 1,\pm 2$.
+Therefore, mhs theory describes particles with helicities $0,\pm 1,\pm 2$, encoded in the symmetric part $\tilde{h}_{\mu\nu}$, and a scalar particle, encoded in the antisymmetric part $\tilde{b}_{\mu\nu}$ – this is the expected result, since, as well known, a two-form gauge field in four dimensions is dual to a scalar particle.
+Moreover, the subsector with vanishing antisymmetric part is the same as in covariant fracton gauge theory, when the special case $\frac{\beta}{\alpha}=2$ is excluded, since, using the identification of the constants in (7.8), $\frac{\tilde{\alpha}_{1}+\alpha_{3}}{2\alpha_{3}}=\frac{2\alpha+\beta}{2\beta}$, so that the first solution is the same as the first solution in (6.9).
+Therefore, the statement in (7.9) holds also in the stronger version
+$\mathscr{M}_{\text{fr}}(\tfrac{\beta}{\alpha}\neq 2)={\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}}_{\big|_{b=0}}.$ (7.11)
 8 BRST formulation
-In this section we formulate the brst rules (2.1) defining the covariant fracton gauge theory in an equivalent way, showing in which sense the covariant fracton gauge theory can be deduced from linearised mhs theory.
-Introduce two doublets $(B_{\mu\nu},\tilde{b}_{\mu\nu})$ and $(r_{\mu},\tilde{\xi}_{\mu})$, with the following transformations
+In this concluding section brst rules (2.1) defining covariant fracton gauge theory are formulated in an equivalent way, off-shell realising the embedding of covariant fractons in linearised mhs theory at the level of symmetries.
+Consider two doublets $(B_{\mu\nu},\tilde{b}_{\mu\nu})$ and $(r_{\mu},\tilde{\xi}_{\mu})$, with the following transformations
 $s\,B_{\mu\nu}=\tilde{b}_{\mu\nu},\quad s\,\tilde{b}_{\mu\nu}=0,$ (8.1a)
 $s\,r_{\mu}=\tilde{\xi}_{\mu},\quad s\,\tilde{\xi}_{\mu}=0,$ (8.1b)
-requiring that both $\tilde{b}_{\mu\nu}$ and $r_{\mu}$ are commuting with ghost number zero, so that $B_{\mu\nu}$ is anticommuting with ghost number $-1$ and $\tilde{\xi}_{\mu}$ is anticommuting with ghost number one.
-The local $s$-cohomology on the jet space $\{h^{(\text{fr})}_{\mu\nu},\lambda\}$ is equivalent to the local $s$-cohomology on the enlarged space including the two doublets $\{h^{(\text{fr})}_{\mu\nu},\lambda,B_{\mu\nu},\tilde{b}_{\mu\nu},r_{\mu},\tilde{\xi}_{\mu}\}$.
+requiring that both $\tilde{b}_{\mu\nu}$ and $r_{\mu}$ are commuting with ghost number zero, so that $B_{\mu\nu}$ is anticommuting with ghost number 1-1 and $\tilde{\xi}_{\mu}$ is anticommuting with ghost number one.
+Then, using the doublet theorem Brandt:1989rd ; Piguet:1995er , the local $s$-cohomology on the space $\{h^{(\text{fr})}_{\mu\nu},\lambda\}$ and derivatives is equivalent to the local $s$-cohomology on the enlarged space including the two doublets $\{h^{(\text{fr})}_{\mu\nu},\lambda,B_{\mu\nu},\tilde{b}_{\mu\nu},r_{\mu},\tilde{\xi}_{\mu}\}$.
 If we redefine $\tilde{b}_{\mu\nu}$ and $\tilde{\xi}_{\mu}$ in this way
@@
 for some $b_{\mu\nu}$ and $\mathring{\xi}_{\mu}$, the last one being anticommuting, then the consistency of the brst transformations implies that
-$s\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},\quad s\,\mathring{\xi}_{\mu}=0,$ (7.3)
+$s\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},\quad s\,\mathring{\xi}_{\mu}=0.$ (8.3)
 Notice that
@@
 The local $s$-cohomology on $\{h_{\mu\nu}^{\text{(fr)}},\lambda\}$ is equivalent to that on $\{B_{\mu\nu},h_{\mu\nu},b_{\mu\nu},r_{\mu},\mathring{\xi}_{\mu},\lambda\}$.
-In other words, the doublet $(r_{\mu},\tilde{\xi}_{\mu})$ implements the condition “$\mathring{\xi}_{\mu}=\tfrac{1}{2}\,\partial_{\mu}\,\lambda$”, which is implied by “$r_{\mu}=0$”; and the doublet $(B_{\mu\nu},\tilde{b}_{\mu\nu})$ implements the condition “$b_{\mu\nu}=\partial_{\mu}\,r_{\nu}-\partial_{\nu}\,r_{\mu}$”, which is implied by “$B_{\mu\nu}=0$”.
+In other words, the doublet $(r_{\mu},\tilde{\xi}_{\mu})$ implements the condition $\mathring{\xi}_{\mu}=\tfrac{1}{2}\,\partial_{\mu}\,\lambda$”, which is implied by $r_{\mu}=0$”; and the doublet $(B_{\mu\nu},\tilde{b}_{\mu\nu})$ implements the condition $b_{\mu\nu}=\partial_{\mu}\,r_{\nu}-\partial_{\nu}\,r_{\mu}$”, which is implied by $B_{\mu\nu}=0$”.
 In this sense, the covariant fracton gauge transformations can be thought as the restriction of linearised diffeomorphisms to the longitudinal ones, as they were firstly introduced Blasi:2022mbl , and covariant fracton gauge theory can be thought as the restriction of the linearised mhs theory in the case in which the antisymmetric part of the linearised vielbein is $\mathrm{d}$-exact.
@@
 $\mathring{t}_{\mu\nu\varrho}=s\,(\tfrac{1}{6}\,\partial_{[\mu}\,B_{\varrho\nu]}).$ (8.9)
-This is the reason why the constants $\alpha_{1}$ and $\alpha_{2}$ appear only in the combination $2\,\alpha_{1}+\alpha_{2}$ n embedding covariant fracton gauge theory in linearised mhs theory.
+This is the reason why the constants $\alpha_{1}$ and $\alpha_{2}$ appear only in the combination $2\,\alpha_{1}+\alpha_{2}$ in embedding covariant fracton gauge theory in linearised mhs theory.
 Acknowledgments
-The author is deeply indebted with F. Fecit and C. Imbimbo for encouragements and illuminating discussions and observations.
+The author is grateful to F.
+Fecit for encouragements and discussions, and he is deeply indebted with C.
+Imbimbo, who read the first draft of this work, pointing out a mistake and suggesting how improve it, stimulating the investigation in Sections 6 and 7.
 Appendix A Weitzenböck geometry within Cartan formalism
-In the formalism introduced by Cartan, one considers the vielbein one-form $e^{a}=e_{\mu}{}^{a}\,\mathrm{d}x^{\mu}$ and the spin connection one-form $\omega^{a}{}_{b}=\omega_{\mu}{}^{a}{}_{b}\,\mathrm{d}x^{\mu}$, and define the torsion two-form $T^{a}=\frac{1}{2}\,T_{\mu}{}^{a}{}_{\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ and the curvature two-form $R^{a}{}_{b}=\frac{1}{2}\,R_{\mu\nu}{}^{a}{}_{b}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ in the following way:
+In the formalism introduced by Cartan, one considers the vielbein one-form $e^{a}=e_{\mu}{}^{a}\,\mathrm{d}x^{\mu}$ and the spin connection one-form $\omega^{a}{}_{b}=\omega_{\mu}{}^{a}{}_{b}\,\mathrm{d}x^{\mu}$, and defines the torsion two-form $T^{a}=\frac{1}{2}\,T_{\mu}{}^{a}{}_{\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ and the curvature two-form $R^{a}{}_{b}=\frac{1}{2}\,R_{\mu\nu}{}^{a}{}_{b}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ in the following way:
 $\ignorespaces T^{a}(e,\omega)$ $=\mathrm{d}\,e^{a}+\omega^{a}{}_{b}\,e^{b},$ (A.1a)
-$R^{a}{}_{b}(\omega)$ $=\mathrm{d}\,\omega^{a}{}_{b}+\omega^{a}{}_{c}\,\omega^{c}{}_{b}.$ (A.1b)
+[…]
 $\omega_{\mu}{}^{a}{}_{b}$ serves as a frame connection (for flat indices).
 The basis or affine connection $\Gamma_{\mu}{}^{\varrho}{}_{\nu}$ (for curved indices) is related to $\omega_{\mu}{}^{a}{}_{b}$ by the requirement that the full covariant derivative of the vielbein, both with respect to $\omega_{\mu}{}^{a}{}_{b}$ and with respect to $\Gamma_{\mu}{}^{\varrho}{}_{\nu}$ (“second vielbein postulate”), vanishes [^16] This is equivalent to define the covariant derivative with respect to the affine connection $\nabla_{\mu}(\Gamma)$ in terms of the covariant derivative with respect to the spin connection $\nabla_{\mu}(\omega)$ in this way: $\nabla_{\mu}(\Gamma)\,(e^{\varrho}{}_{a}\,V^{a})=e^{\varrho}{}_{a}\,\nabla_{\mu}(\omega)\,V^{a}$, for any $V^{a}$.
-Indeed on a side $\nabla_{\mu}(\Gamma)\,(e^{\varrho}{}_{a}\,V^{a})=\partial_{\mu}\,V^{\varrho}+\Gamma_{\mu}{}^{\varrho}{}_{\nu}\,V^{\nu}$; on the other $e^{\varrho}{}_{a}\,\nabla_{\mu}(\omega)\,V^{a}=\partial_{\mu}\,V^{\varrho}+e^{\varrho}{}_{a}\,(\partial_{\mu}\,e_{\nu}{}^{a}+\omega_{\mu}{}^{a}{}_{b}\,e_{\nu}{}^{b})\,V^{\nu}$, so that we obtain (A.2).
+Indeed on one side $\nabla_{\mu}(\Gamma)\,(e^{\varrho}{}_{a}\,V^{a})=\partial_{\mu}\,V^{\varrho}+\Gamma_{\mu}{}^{\varrho}{}_{\nu}\,V^{\nu}$; on the other $e^{\varrho}{}_{a}\,\nabla_{\mu}(\omega)\,V^{a}=\partial_{\mu}\,V^{\varrho}+e^{\varrho}{}_{a}\,(\partial_{\mu}\,e_{\nu}{}^{a}+\omega_{\mu}{}^{a}{}_{b}\,e_{\nu}{}^{b})\,V^{\nu}$, so that we obtain (A.2).
 VanNieuwenhuizen:1981ae :
@@
 if the vielbein is invertible. If we define
-$\ignorespaces T_{\mu}{}^{\varrho}{}_{\nu}(\Gamma)$ $=\Gamma_{[\mu}{}^{\varrho}{}_{\nu]},$ (A.4a)
-$R^{\varrho}{}_{\sigma\mu\nu}(\Gamma)$ $=\partial_{[\mu}\,\Gamma_{\nu]}{}^{\varrho}{}_{\sigma}+\Gamma_{[\mu}{}^{\varrho}{}_{\alpha}\,\Gamma_{\nu]}{}^{\alpha}{}_{\sigma},$ (A.4b)
+[…]
+[…]
 then they are related to the Cartan definitions (A.1a)–(A.1b) according to
-$\ignorespaces T_{\mu}{}^{\varrho}{}_{\nu}(\Gamma(e,\omega))$ $=T_{\mu}{}^{a}{}_{\nu}(e,\omega)\,e^{\varrho}{}_{a},$ (A.5a)
-$R^{\varrho}{}_{\sigma\mu\nu}(\Gamma(e,\omega))$ $=R_{\mu\nu}{}^{a}{}_{b}(\omega)\,e^{\varrho}{}_{a}\,e_{\sigma}{}^{b}.$ (A.5b)
+[…]
+[…]
 This follows straightforwardly by replacing (A.3) in (A.4a)–(A.4b).
@@
 The corresponding affine connection, according to (A.3), is again a Cartan-Maurer form: [^17] The author thanks C. Imbimbo for pointing this out.
-$\Gamma_{\mu}{}^{\varrho}{}_{\nu}(e,\omega^{\text{(flat)}})=e^{\varrho}{}_{a}\,\partial_{\mu}\,e_{\nu}{}^{a}+{\omega^{(\text{flat})}}_{\mu}{}^{a}{}_{b}\,e^{\varrho}{}_{a}\,e_{\nu}{}^{b}=(\lambda_{c}{}^{b}\,e^{\varrho}{}_{b})\,\partial_{\mu}\,(\lambda^{c}{}_{a}\,e_{\nu}{}^{a})$ (A.10)
+[…]
 reflecting the fact that it is flat: $R(\Gamma(e,\omega^{\text{(flat)}})=0$.
@@
 This is called Weitzenböck condition, which breaks the local Lorentz invariance, but it preserves the global Lorentz invariance.
-A useful tensor is the contorsion, which is defined as the difference between an arbitrary affine connection and the Levi-Civita connection [^11] Using (A.3), we can also write the contorsion in terms of an arbitrary spin connection and the torsionless one: $-K_{\mu}{}^{\varrho}{}_{\nu}=(\omega_{\mu}{}^{a}{}_{b}-{\omega^{\textsc{(gr)}}}_{\mu}{}^{a}{}_{b})\,e^{\varrho}{}_{a}\,e_{\nu}{}^{b}$.
-$-K_{\mu}{}^{\varrho}{}_{\nu}(\Gamma):=\Gamma_{\mu}{}^{\varrho}{}_{\nu}-{\Gamma^{(\textsc{gr})}}_{\mu}{}^{\varrho}{}_{\nu},$ (A.13)
+A useful tensor is the contorsion, which is defined as the difference between an arbitrary affine connection and the Levi-Civita connection [^18] Using (A.3), we can also write the contorsion in terms of an arbitrary spin connection and the torsionless one: $K_{\mu}{}^{\varrho}{}_{\nu}=(\omega_{\mu}{}^{a}{}_{b}-{\omega^{\textsc{(gr)}}}_{\mu}{}^{a}{}_{b})\,e^{\varrho}{}_{a}\,e_{\nu}{}^{b}$.
+$K_{\mu}{}^{\varrho}{}_{\nu}(\Gamma):=\Gamma_{\mu}{}^{\varrho}{}_{\nu}-{\Gamma^{(\textsc{gr})}}_{\mu}{}^{\varrho}{}_{\nu}$, (A.13)
 which implies that $K_{\mu\varrho\nu}=K_{\mu}{}^{\sigma}{}_{\nu}\,g_{\varrho\sigma}$ is antisymmetric in the last two indices.
 Notice that the antisymmetric part in the first and in the last index is the torsion (A.5a):
-$-K_{[\mu}{}^{\varrho}{}_{\nu]}(\Gamma)=T_{\mu}{}^{\varrho}{}_{\nu}(\Gamma).$ (A.14)
+[…]
 This expression can be solved for writing the contorsion in terms of the torsion
@@
 $=-\tfrac{1}{2}\,(T_{\mu\varrho\nu}+T_{\varrho\mu\nu}+T_{\varrho\nu\mu})=-\tfrac{1}{2}\,(T_{\mu[\varrho\nu]}+T_{\varrho\mu\nu}).$ (A.15)
-We can relate the Levi-Civita connection, the Weitzenböck connection and the Weitzenböck torsion in the following way. On a side, using (A.13),
-$-K_{\mu}{}^{\varrho}{}_{\nu}(W)=W_{\mu}{}^{\varrho}{}_{\nu}-{\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}.$ (A.16)
+We can relate the Levi-Civita connection, the Weitzenböck connection and the Weitzenböck torsion in the following way. On one side, using (A.13),
+$K_{\mu}{}^{\varrho}{}_{\nu}(W)=W_{\mu}{}^{\varrho}{}_{\nu}-{\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}$. (A.16)
 On the other, using (A),
-$-K_{\mu}{}^{\varrho}{}_{\nu}(W)=\tfrac{1}{2}\,(T_{\mu}{}^{\varrho}{}_{\nu}(W)-T_{\mu\nu}{}^{\varrho}(W)+T^{\varrho}{}_{\mu\nu}(W)).$ (A.17)
+$).-K_{\mu}{}^{\varrho}{}_{\nu}(W)=\tfrac{1}{2}\,(T_{\mu}{}^{\varrho}{}_{\nu}(W)-T_{\mu\nu}{}^{\varrho}(W)+T^{\varrho}{}_{\mu\nu}(W$)). (A.17)
 Therefore,
-${\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}=W_{\mu}{}^{\varrho}{}_{\nu}+\tfrac{1}{2}\,(T_{\mu}{}^{\varrho}{}_{\nu}(W)-T_{\mu\nu}{}^{\varrho}(W)+T^{\varrho}{}_{\mu\nu}(W)).$ (A.18)
+$).{\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}=W_{\mu}{}^{\varrho}{}_{\nu}+\tfrac{1}{2}\,(T_{\mu}{}^{\varrho}{}_{\nu}(W)-T_{\mu\nu}{}^{\varrho}(W)+T^{\varrho}{}_{\mu\nu}(W$)). (A.18)
 Appendix B Computation of a BRST cohomology
@@
 $H(s)=\{\mathring{\xi}^{\mu},\mathring{T}_{\varrho\nu\mu},\partial_{\alpha_{1}}\,\mathring{T}_{\varrho\nu\mu},\dots,\partial_{\alpha_{n}}\dots\partial_{\alpha_{1}}\,\mathring{T}_{\varrho\nu\mu},\dots\},$ (B.4)
-In order to show this, consider the decomposition $\partial_{\varrho}\,\overline{h}_{\mu\nu}$ in irreducible components.
+In order to show this, consider the decomposition $\partial_{\varrho}\,\overline{h}_{\mu\nu}$ in irreducible components: [^19] See note 2 for the notation of Young tableaux.
 $(1)\otimes[(1)\oplus(2)_{t}\oplus(1,1)]=(1)\oplus[(1)\oplus(2,1)_{t}\oplus(3)_{t}]\oplus[(1)\oplus(2,1)_{t}\oplus(1,1,1)],$ (B.5)
@@
 $t^{\textsc{(hs)}}_{\lambda\mu\nu}$ $=\tfrac{1}{2}\,({T^{\textsc{(hs)}}}_{\lambda\mu\nu}+{T^{\textsc{(hs)}}}_{\mu\lambda\nu})+\tfrac{1}{6}\,(g_{\nu\lambda}\,v_{\mu}+g_{\nu\mu}\,v_{\lambda})-\tfrac{1}{3}\,g_{\lambda\mu}\,v_{\nu}=$
-$=\tfrac{1}{2}\,(T_{\nu\lambda\mu}+T_{\nu\mu\lambda})+\tfrac{1}{6}\,(g_{\nu\lambda}\,T_{\mu\varrho}{}^{\varrho}+g_{\nu\mu}\,T_{\lambda\varrho}{}^{\varrho})-\tfrac{1}{3}\,g_{\lambda\mu}\,T_{\nu\varrho}{}^{\varrho}.$ (C.6d)
+$=\tfrac{1}{2}\,(T_{\nu\lambda\mu}+T_{\nu\mu\lambda})+\tfrac{1}{6}\,(g_{\nu\lambda}\,T_{\mu\varrho}{}^{\varrho}+g_{\nu\mu}\,T_{\lambda\varrho}{}^{\varrho})-\tfrac{1}{3}\,g_{\lambda\mu}\,T_{\nu\varrho}{}^{\varrho}$.
+(C.6d)
 $v_{\mu}$, $a_{\mu}$ and $t_{\lambda\mu\nu}$ are the irreducible components of the torsion in 4d, with $a_{\mu}$ (4d dual to) the totally antisymmetric part.
@@
 $c_{1}=c_{2}=c_{3}=0.$ (C.11)
+Appendix D Polarisation matrices
+See for example VanNieuwenhuizen:1981ae . Choosing the generator of $z$-axis rotation as
+$J_{z}=\begin{pmatrix}0&0&0&0\\ 0&0&-i&0\\ 0&i&0&0\\ 0&0&0&0\end{pmatrix}$, (D.1)
+the polarisation vectors with helicities $\pm 1$ are
+$\varepsilon_{\mu}^{\pm}=\tfrac{1}{\sqrt{2}}\,(0,1,\pm i,0),$ (D.2)
+such that
+$J_{z}\,\varepsilon_{\mu}^{\pm}=\pm\,\varepsilon_{\mu}^{\pm}.$ (D.3)
+Define the polarisation matrices as
+$\ignorespaces\lambda_{\mu\nu}^{0}=\varepsilon_{(\mu}^{+}\,\varepsilon_{\nu)}^{-}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&-i&0\\ 0&i&0&0\\ 0&0&0&0\end{pmatrix}$, (D.4a)
+$\lambda_{\mu\nu}^{\pm 1}=\tfrac{1}{\sqrt{2}}\,p_{(\mu}\,\varepsilon_{\nu)}^{\pm}$ $=\frac{1}{2}\begin{pmatrix}0&-1&\mp i&0\\ -1&0&0&1\\ \mp i&0&0&\pm i\\ 0&1&\pm i&0\end{pmatrix}$, (D.4b)
+$\lambda_{\mu\nu}^{\pm 2}=\varepsilon_{(\mu}^{\pm}\,\varepsilon_{\nu)}^{\pm}$ $=\frac{1}{2}\begin{pmatrix}0&0&0&0\\ 0&1&\pm i&0\\ 0&\pm i&-1&0\\ 0&0&0&0\end{pmatrix}$. (D.4c)
+They are such that
+$J_{z},\lambda_{\mu\nu}^{s}]=s\,\lambda_{\mu\nu}^{s}.$ (D.5)
+Two useful formulas are the following:
+$\begin{pmatrix}0&0&0&0\\ 0&a&b&0\\ 0&b&-a&0\\ 0&0&0&0\end{pmatrix}=(a-i\,b)\,\lambda^{+2}+(a+i\,b)\,\lambda^{-2},\quad\begin{pmatrix}0&-a&-b&0\\ -a&0&0&a\\ -b&0&0&b\\ 0&a&b&0\end{pmatrix}=(a-i\,b)\,\lambda^{+1}+(a+i\,b)\,\lambda^{-1}.$ (D.6)